A Complete Bibliography of Publications in the Journal of the Society for Industrial and Applied Mathematics. Series A, On control and SIAM Journal on Control

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

23 August 2019 Version 2.03

Title word cross-reference

(1 ≤ p ≤ ∞) [WD69]. (Q) [CS74]. 2 × n [SD71]. GL(n, R) [Sus72]. H^{2} [BB75]. l [DP72]. L^{p} [WD69]. L_{2} [Fre68]. L_{p} [tC68]. min sup [DM69]. N [Ben74, KD66, SRP69, SP71a, Sca74b, Var70, Paz67]. n × n [Paz67]. σ [RT73]. τ [LH69]. ε [Roo75a]. W_{2}^{(1)} [BJL75].

-Capture [DP72]. -decomposition [LH69]. -dimensional [Paz67]. -Functions [BB75]. -Person [SP71a, Sca74b, SRP69, Var70]. -Representations [Roo75a]. -stability [tC68, Fre68]. -stage [KD66].

10 [Luk73].

62 [Vin73].

9 [CLS73].

Abelian [FFZ69]. Abstract [Ber74, Jur70, MP71b, Dan73, Neu66, Neu67]. Accelerated [Mey74]. Accuracy [BJ71]. Achieve [Sus75]. adaptive [KP73]. Addendum [Bha66]. Additive [Glo69]. Admissible [Mir69]. Aerodynamics [Mie65]. Affine [Cur75b, DM72b]. Algebraic [AM68, Mol73a, Mol73b]. Algorithm [Bar72a, CC68, CM70b, KP72, Lin74a, McG65, Mer64, Pau75, PH75, Sch73, CE66].

Algorithms [EZ71, Kle74, MP71b, Mey74, TV67]. Allocations [WS74b]. Almost [AJ71]. Amplitude [SR64]. Amplitudes [SL70]. Analysis [Doo64, Hol68, Mic74, Pap73, SG63, Kir67, LH69, NC68, Sze62]. Analytic [CL75, Cul64, Jur72]. Analyticity [FFA69]. Antizigzagging [Kle74]. Aperiodic [JP64]. Appearing [HH63, Lo75]. Application [Bai65, Kai69, Rob75, SP71b, TAG70, CE66, NY68, Rus69b]. Applications [BVW75b, Cur75b, For75, Kir67, Kus64, Pet72, Won71, Won64, Gam65, Neu66, Neu67, Onu71, Por68, Neu67]. Applied [Smi64]. Approach [Bal72a, Bal72b, Kel64, Pap73, WM70, DIR73, Hay66]. approximate [Bri67, Bri68]. approximating [SS71]. Approximation [Bal74, Bal75, Ste73, de 68, Jac69a]. Approximations [BJ71, BBS69, Cul72, Cul69]. Arbitrary [Gur64, MF71]. Arc [McC75]. arising [Ben66]. Ascending [EH64]. Aspects [Lei65]. Assignment [MF71, MW70, WM70]. associated [Zak67]. Asymptotes [Kra64a]. Asymptotic [BD68, BB64, Buc66, San75, Ste72, Zlo70]. attainable [Her67, Her68]. Augmented [Roc74]. Automata [Arb65, GS66, de 68, CS70]. Autonomes [Fau72]. Autonomous [Bra72, Chu74, Dat74].

Backward [Lin74b]. Balakrishnan [JM70, Vin74]. ballistics [Has68]. Banach [Bal65, Dat70, Gui69, KR66, Mey70a, Tri75, Vin74, Zlo70]. Band [Bal74, Bal75]. Band-Limited [Bal74, Bal75]. Bang [Hal64, Kre74, SV64, Sus72, Aro73, Fat68b]. bang-bang [Aro73]. based [Ben70]. Bases [For75]. beam [Kom68]. Behavior [Ste72]. Bellman [Gir69]. Best [Rom69]. Between [Gir69, Pea66b]. Bilinear [DIR74, Kre75, RM68]. Bilinearization [Lo75]. binary [Paz67]. Bolza [Bar75a, Bar75b, Goh66b, Roc72]. Bound [CC68, PD70]. Boundary [Cir69, Fat68a, Fat75, Fly73, GR75, Mer64, Rus71a, Rus71b, Wan75, PP73, Rus73]. Bounded [Gui65a, MN72, Pen72, Rus64, SL70, Tri75, VL66, For67, Had73]. bounded-impulse [For67]. Bounded-input [VL66, Had73]. bounded-output [Had73, VL66]. Boundedness [AJ71, Zak67]. Bounds [BJ71, Rob75, Wit70]. Broad [Vid72, Neu66, Neu67].

Calculus [War65a]. Canonical

[Aka75, DIR73]. capable [Ris66]. Capture [DP72]. Cas [Ben74]. Case [Cul64, DP72, Fly74a]. Cauchy [AT75]. Causal [Dol73, FF69]. Causality [De 74, FFA69, Wil69, Wit71, DeS73]. Certain [AY70, Bar72b, Ces75, Gir69, Jur72, MN72, Mis65, Ole74, Sus72, Aro73, Bru68, Ego66, Fil62, Kul66, Kul67]. Certaines [Fau72]. chain [Kul66, Kul67]. Chains [MS75]. characteristics [Fre68, Joh73, RT73]. Characterization [Aub70, BJL75, Ole74, WE75]. Chow [Kre74]. Circle [Coo75]. Class [Bar71, Cul64, Mer64, Mur74, SP71a, Vid72, BD68, Bal63, Ben70, Ber67, CE66, HK66, NN66, NY68, Neu66, Neu67, SRP69, Sno64]. Classes [Roo75a, de 68, LL66]. Classical [Fil67]. classification [CP66]. Closed [Bru74, Kre69, Cul66a]. Closed-Loop [Bru74]. Closest [MDM74]. Closure [Ber74, Ces71, Ole74, Bru68, Her67, Her68]. Coefficients [San64, Leh67]. coinciding [HO73]. Combined [Ber75]. Comments [NW70, Wei77, Vin73]. Common [Arb65]. Compact [Lob74, FFZ69]. Comparison [Rao72]. Compensation [MW70]. Complete [Chu65, Fat66, Tro71, Sle74]. Completely [RM68]. Complex [ABI71, Leh67]. Components [BT64].

Comportement [Fau72]. Computational [CL75, Bar69]. Computing [AS65, Bal68, HB63, Gil66]. Concept [IS75, Hac73]. concerning [Ris66]. Condition [Lee63, WS74b, Bru68, Hac73, Jac69b]. Conditional [Kus64]. Conditions [ABI71, AT75, Aub70, Bar71, BC72, Ces74, Dav72b, FG70, GK72, Geh74, Jac70, JM70, Kus65b, Kus72, MT72, MP71a, MP69, MN72, Ros64, Rus75, SP71a, Sko74, Vır70, WW69, Wan75, Zol72, Ban68, Ber67, Bol66, DV73, DK70, Ego67, Goh66a, Gui69, Hau73b, How73, Man66, Moy73, Ris70, SRP69, Sto73, Sur73, ZF68, Zlo70]. Conflict [Gin67a]. Conjugate [Haa75, KP72]. Connected [Mis65]. Connections [ADT75]. considerations [Mae73]. Considered [Gui65b]. Consistency [NS75]. constant [Leh67]. Constrained [Ber75, Bra74, CCP66, Cul72, Dan71, Kle74, WS74a]. Constraint [Fly73, Cul66a, Hau73b]. Constraints [Aub70, Bis74, Blu65, CH75, FG70, Geh74, HB63, JM70, MN74, McG65, Nor71, Roc72, Rub72, Rus75, Sas74, Sca74b, Sch75b, Sch75c, TV75, Vır70, Mor66, PP73, Vac66, Wah73]. Construction [ABL75, For67, LwL67, RT73, Fre68]. Consumption [Pel72]. Contain [Ris65]. Containing [Sch65]. Continuity [Rus69b]. Continuous [Chi69, Cul72, Hal74, Kak72, Kus72, Lin74b, Mau75, Pli75, ARS67, Cul69, Gri67, Mil68, Ris70, Wah73]. Continuous-Time [Lin74b]. Contrôlabilité [Lob70, Lob76]. contraction [OJ68]. Control [AT74, Arb65, Arm74, Art75, Bal65, Bal68, Bal72b, BK72, Bar75a, Bar75b, Bar71, Bar72a, Bar72b, Bar73, Bau72, BB64, Ben71, BV75, Ber74, Bic68, Bis74, BSB63, BT64, Bru74, Buc66, BBS69, Ces65a, CLS73, CS74, Chi69, Chu74, Chu75, Cul66b, Cul72, Cul64, Dat73, Dau71, DM72b, Dem68, Dre64, DV71, DV75, Fat64, Fat68a, Fat75, Fle68, Fle71, FG70, Gal69, GK74, Gin67a, Gin67b, Gol65a, GSG67, GR75, Háj71, Har64, Hau71, HH63, Her72, Hes65, HB63, Hol74, Hol75, Hua72, Jac70, JK73, JP64, Jur70, Jur72, Kel64, Kom70, Kre69, Kre73, Kre74, Kus65b, KB70, LaS62, LwL67, Lee63, LCB72, Lin73, Lo75, LW75, Luk71, Luk72a, Luk73, MN74, Mal69, Mau75, McG65, Mey70a, MN72, Mir69, MH68, MS75, Neu62, Nis67]. Control [Nor67, Oğu63a, Oğu63b, Ogu64, Oka64, Pea66b, Pon68, Por66, Psh66, Ris65, Roc72, Roo69, Rox65, Rus64, Rus71a, Rus71b, Sas72, SR64, Sch65, Sch75c, SL70, SV64, Sus72, Swo64, SG63, TAG70, Wan75, War70a, War70b, Wit68, Won68a, Won68b, Zab74, Zab75, Zmo74, Bal63, Ban68, Ben66, Ber67, Bru68, BEL68, CE66, Ces71, CP66, Cul66a, Cul69, Dan73, Dan66, Dau73, Dav73, Fat68b, Fil62, Gam65, Gil63, Goh66a, Hau73a, HO73, Hay66, Hol68, Jac67b, Jac69b, Joh73, Kir63, KP73, Kom68, Kus67b, Kus68, Man66, MB73, Moy73, NY68, NW70, Nor68, PP73, Por68, Ris70, Ros66, RFL69, Rub66, Rus69b, Rus73, SS71, Sch67, Sle74, Sto73, Sze62, Vac66, VW66, Vin69a, Vin69b, Vin69c, Vin70, Vin73, Won67, Won69]. Control-Phase [MN74]. Controllability [Bra72, Che70, Chu65, Chu75, Cir69, Con67, DM72a, DM72b, Fat66, Gil63, HH70, Her64, Her74, Jur70, Jur72, LS75, Lob70, Lob74, Lob76, Luk72b, Luk74, Mar65, MF71, Mor71, Sak74, SY71, Sch75b, SM67, SA68, Sus75, Tri75, Tro71, Tsu70, WE75, Wei67, Wei72, Yor72, Zmo74, Aro73, DK70, FJ68, For67, Luk73, Sap73, Sle74]. Controllable [Pop72, RM68]. Controlled [BJL75, Pli73]. Controllers [Bra72]. Controls [AKR67, Bau72, Bri63, Bri64, CLS71, Hal68b, Kus65a, Lef65, McS67, SR64, Sch72, War65b, War67a, War67b, War74, Zol72, AD66, Bri67, Bri68, Ces68, CLS73, DS71, For67, Gam63, Gri67, Hal71, LL66, Rus66, SY71, Sap73, Sno64]. Convergence [AY70, Art75, CC68, Dan71, Kle74, SS71, TV67]. Convergent [Bar72a]. Converse [Kus67a]. Converses [Gor72a]. Convex [Abr75, Bar75a, Gol65a, GT73, Man65, PN70, Rob75, Roc72, DR67, Gil66]. Convexity [CS74, Ces74, Ces75, Yor71, Bru68, Her67, Her68, HH66]. Convolution [DC72]. Coordinate [MN72, Rus64]. Coordinates [Gui65a, Hua72, Sca74a, Paz67]. Coprime [Vid75]. Correction [KH69, Wei77]. correctness [Kir63]. correlated [PD70]. Coset [Bro72]. Cost [Chy67, Dat73, Dat74, DMM75, KB70, TV75, BM73]. Costly [Rus69a]. Costs [Luk71, Luk72a]. Countable [Kak72]. Counterexample [Wit68]. covariance [PD70]. Coverability [WS74b]. Criteria [Ack75, Coo75, Dav72a, Haa75, Kra64b, Man63, Nol75, Sko74, Ven71, FZ68, Rus73]. Criterion [And67, Lel70, LR69, Zak69, Had73]. Curvature [CH75, Rub72]. Curves [Rub75, War65a, Gam63]. Cutting [EZ71]. Cycles [Fau72].

Dans [Ben74]. Data [AS65, ABL75, Hal74, Pap73]. Davidon [TAG70]. Decision [Kak72, Pli75, Ben70, CS70, Mil68]. Decomposition [MA74, Pea66a, Var66, DIR73, DeS73, LH69, MA74]. Decoupling [FW74, MW70, SP71b, WM70, Gil69]. deducing [Bal67]. Defined [MF71, War70b]. definite [CS70]. Degree [LW75]. Delay [Hal68a, Oğu63a, Wei67, Zmo74, DS71]. Delay-Differential [Wei67, DS71]. Delays [CL66, Wan75]. Densities [Kus64]. Dependent [Hau71, Ros64, Hau73a, Won67]. Depending [BT75]. derivation [Kul66, Kul67]. Derived [Dav72b]. Descent [Gol65b, Bal63, BM73, Vac66]. Described [Gal69, Oğu63b, Hal66, NW70, Vin69a, Vin69b, Vin69c, Vin70, Vin73]. Describing [BF71, Dol73]. Description [Kal63, Pop72]. Design [Hof75]. Detectability [Hew75]. detection [Par62]. Deterministic [IMK75, MS75]. Difference [BBS69, Gor72a, Gor72b, Oğu63b, Hal66, RFL69]. Different [Pap73]. Differentiability [Hal74, MT72]. Differential [AT75, Ban69, BK72, Bha64, Cas69, Cur75b, DM72b, DMM75, EKM73, Fat64, GK74, Hal68a, Kus64, KB70, Kwa75, Man63, Mil72, Mir69, Mis65, Mis74, Nol75, Oğu63b, PR75, Pau75, Pon65, Pon74, Rao72, Ris71a, Ros64, Sak74, Sak75, San64, SP71a, Sca74a, Sca74b, Sch65, SH71, Ste72, Ste73, Ste74, SG63, Var67, VL69, Wei67, Won64, BJ70, Ber67, Bha66, CM70a, DS71, Fil67, Gam65, NN66, Onu71, RFL69, Rus69b, SRP69, Sur73, Sze62, VL66, Var70, Wei68, Zak67]. differential-difference [RFL69]. Différentielles [Fau72]. Differentiels [Ben74]. Diffusion [GSG67, Kus67b]. Diffusions [AT74, Fle68, Pli73]. Dimensional [BB75, Bau72, Cul72, Cur75a, Cur75b, DM72a, Rus71a, Rus71b, SB74, CCP66, Fat68b, Paz67, Rus66, ARS67]. Dimensions [WE75, Hay66, Her67, Her68]. direct [Leh67]. direction [TV67]. Directions [Mey73, Sus75]. discontinuous [Mae73]. discounted [Kus67b]. Discrete [Cul69, Dav72b, JP64, LCB72, Lin74a, MS71, MF71, Swo64, Wei72, Zab74, Zab75, Con67, HH66, Med67, NC68]. Discrete-Time [JP64, LCB72, Lin74a, Wei72, Zab74, Zab75]. Discretional [HH66]. Discretization [Dan71]. Distributed [Bar71, BV75, Kus70, LCB72, LR69, Sas72, Vid75, Zab74, BEL68, Ego66, Ego67, SS71]. distributed-parameter [Ego67]. Distribution [SD71, Prè66]. Distributions [Fat75, Zak69]. Distributive [Por66]. Disturbance [Cul64]. Disturbance-Free [Cul64]. Disturbances [Ris75]. Disturbed [Luk74]. d’Optimisation [Lor73a]. d’optimization [Lor73b]. Doubly [Bre65]. Drivable [Mey73]. driving [PD70]. Dual [Bar72b, Bar73, Ber75, Vin74, PP73]. Duality [MH68, Pal66, Pea66b, Pea66a, Roc74]. d’une [MA74]. Duoplex [Kün66]. Duration [Rao72]. Dynamic [Bel66, DV73, Gir69, MW70, Pal66, Ris71a, Ris75, Bol66, Ris70]. Dynamical [DIR74, De 70, HT70, Hey72, Kal63, Luk69, Sko72, Con67, DIR73, FJ68, LH69]. dynamics [RFL69, Var70, Mor66].

economics [NY68]. effects [Wil71b]. Efficiency [Pel72]. Efficient [KP72, PH75, Bar69]. Effort [BT64, Neu62, Sno64]. Elements [Sch75a]. Elliptic [Zol72]. Encirclements [Dav72a]. Encounter [Mis74]. Energy [Cul64, Joh73]. Equality [Blu65, Geh74, MN74, Nor71, Sch75b, Sch75c]. Equation [Buc75, Buc77, Bur65, Cur75b, GR75, Hew75, KB70, Ris71a, Rus71a, Rus71b, SH71, SG63, TV75, Won68a, Kus68, Mae73, Mol73a, Mol73b, RFL69, Sze62, Won69]. Equations [Ban69, BK72, CL75, Dat74, Dol73, Fat64, Gal69, Gor72a, Gor72b, Hua72, Kus64, Luk74, Man63, Mil72, Mis65, Oğu63b, RBR74, Ros64, Sak74, Sak75, San64, Sch65, Tsu70, War70b, Won64, Yor72, BJ70, CM70a, Cow73, DS71, Fil67, Gam65, Hal66, NN66, NW70, Onu71, Rus69b, Str66, Sur73, Vin69a, Vin69b, Vin69c, Vin70, Vin73, Zak67, Fau72]. Equilibrium [Luk71, Luk72a, PR75, Pau75]. Equivalence [CP66, HT70, Kre73, Mol73a, LL66]. Errata [Bal75, Bri68, Dau73, Gar64, Her68]. Erratum [Buc77, CLS73, Hal71, Lob76, Luk72a, Luk73, Nor68, Vin70, Won69, Lob76]. Error [BJ71, Rob75, Sny75, PD70]. espace [MA74]. Estimates [Fre72, GT73, NS75]. Estimation [Bal72a, IMK75, Kai72, KH69, Pea66b, Sny75, Wei77, PD70]. Estimators [AS65]. Euclidean [Zmo74]. Evading [Mis74]. Evasion [EF75, GK74, Ho66, Por67, Sak70]. Even [Mau75]. Evolution [Tsu70]. Exact [Wol72]. exactness [How73]. Example [Bis74, Mau75, CM70a, Ris66]. Exceptional [CLS71, CLS73]. Exhaustive [LL66]. Existence [AT74, Bau72, Ben70, Ben71, Ber74, Bri63, Bri64, Ces65b, Ces65a, Ces68, CLS71, Ces74, EG74, Fau72, Has68, Hau74, Kus65a, Mey65, MS75, Nis67, PR75, Var67, VL69, WS74a, Zak69, AD66, CLS73, Cow73, Dav73, Gam63, Gri67, Jac67b]. expansion [Paz67]. Expansions [San75]. Expected [Pli75]. Exponential [Bha64, Mur74, Bha66]. Expressions [Sny75]. Extended [Ris65]. Extension [Arm74, Hay66, Kul67]. Extensions [Hac73, Jac67a, Kro68]. exterior [Has68]. Extraction [Par62]. Extremal [AKR67, Mie65, Rub75, Gam65]. Extremals [Haa75, KO75, Goh66a].

Factorizations [Vid75]. Family

[Mey70b, Wah73]. Feasible [Mey73, TV67]. Feedback [Bri63, Bri64, Dav72b, DM72b, DC72, FG71, IMK75, Kra64b, Lin73, Luk71, Luk72a, Sko72, Sko74, SB74, Vid75, Wit71, Wol72, BD68, tC68, Gil69, WD69]. Fibonaccian [Kro68]. Field [Gur64]. Fields [LS75, MF71]. Filippov [Jac67a]. Filter [Rom69, Smi64]. Filtering [Cur75a, Kus70, Lin74a, Lin74b, Won64]. final [PP73]. Finding [MDM74]. Finite [ABL75, Chu75, Cul72, GS66, Glo69, Har64, Mil68, CCP66, Her67, Her68, Paz67, Rus66]. Finite-Dimensional [Cul72, CCP66]. Finite-Turn [GS66]. first [Joh73]. Fixed [Mey70a, Rao72, Gil67, Rub66]. fixed-point [Gil67, Rub66]. fonction [MA74]. Fonctionnelles [Ben74]. Form [PN70, Gil66]. formulation [Bal63, Kir63]. Fourth [SH71]. Framework [Arb65].

Frank [CC68, Mey74]. Fredholm [Dav72a]. Free [Cul64]. Freedom [LW75]. French [Lob70, Lor73b, MA74, Lob76]. Frequency [Lel70]. Fritz [Geh74]. Fuel [Neu65, Gri67]. Function [Bal72b, BK72, BF71, Dan71, EG74, FG71, Fre72, Mey65, Por67, How73, MA74]. Functional [Ban69, BK72, Dat74, KB70, Mol75, Nol75, War70a, BJ70, DR67, Kir67, Wah73]. Functional-Differential [KB70]. Functionals [Blu65, Chy67, DR68, Gol66, SB74, WS74a, Wie72, BM73]. Functions [Art75, BB75, Bel62, Doo64, Gir69, Hal74, Man65, Mur74, Ole74, Pon68, Roc74, Rus64, War67a, War67b, Wil71a, Jac67a, Jac69a, Kus67a]. Further [Bla66, Kro68, Vin73].

gains [FZ68]. Galerkin [Dan73, MB73]. Game [GK74, SD71, Var67]. Games [Cas69, DP72, EKM73, EF75, Hal68a, Luk71, Luk72a, Mir69, Mis74, PR75, Pau75, Pet72, Pon65, Pon74, Por67, Rao72, Sak70, SP71a, Sca74a, Sca74b, Ste72, Ste73, Ste74, VL69, Ber67, SRP69, Var70, Wil71b]. Gaussian [Hol75, IT75, Roo69]. General [Cha66, Fly74a, Kle74, Neu65, Neu66]. Generalization [Hal64, JM70, Kre74, Vin74]. Generalized [Bur65, EZ71, EF75, FF69, Gam63, Gui69, IT75, KH69, McG65, Nor71, Rub75, Wei77, AM68, Bar69]. generating [Paz67]. Generation [Wil71a]. Generic [FW74]. Geometric [Háj71, WM70]. Geometrical [Lei65]. Geometrically [Bar72a, Bra74]. Geometry [Her72, Bla66]. Global [Aro73, Her74, Lo75, Luk72b, Luk73, Luk74, Sap73, Ven71, DK70]. Goddard [Gar63, Gar64]. governed [Kus68]. Gradient [KP72, MT72, KP73]. Gravitational [Gur64]. Green [Hay66]. Group [Bro72, Glo69]. groups [FFZ69]. Guidance [Bre65, Smi64]. hand [Fil67]. Hereditary [BJL75, Cur75b, DM72b, DMM75]. Hestenes [Gui65a, Gui65b]. High [BJ71, GK72]. Higher [Rus71a, Rus71b, Hay66]. Higher-Dimensional [Rus71a, Rus71b]. Hilbert [Ack75, Bar75a, Bar75b, Coo75, De 74, Dol74, DZ75, FFZ69, Par62, Sle74, TAG70, Tsu70, Zab75, Zem74]. Hold [Aub70]. horizon [Mil68]. Hyperbolic [Cir69, Joh73, Rus66, Rus73, Sch67, Sur73]. Hypotheses [Gui65b].

Identification [AY70, NS75, Ris71b, Roo75a, Roo75b, Bel66, OJ68]. Identified [CR68]. identity [Kai69]. II [BVW75b, DZ75, DV75, Lor73b, Neu67, OK75, Roo75b, Rus71b, Vin69b]. II. [Bri64]. III [Vin69c]. Implementations [KP72]. Implicit [Hal74, Jac67a]. improved [RT73]. impulse [For67]. impulsive [Ris66]. Inaccessible [Roo69]. Incompletely [CR68]. independent [Sch67]. Index [MS73, AD66, Sak66]. Index-Invariant [MS73]. Induced [ADT75]. inequalities [HK66]. Inequality [HB63, JM70, MN74, McG65, Rus75, Vac66, Wah73]. Infima [Abr75]. Infinite [BB75, Cur75a, Cur75b, DM72a, DMM75, SB74, Fat68b, Her67, Her68]. Infinite-Dimensional [BB75, Cur75a, DM72a, SB74, Fat68b]. Infinite-Time [DMM75]. Information [Wit71, Ben70, Kul66, Kul67]. information-theoretic [Kul66, Kul67]. Initial [JM70, PP73]. Innovation [Lin74b]. Innovations [Kai72]. Input [ABL75, Kre75, SP71b, Vid72, Wil71a, Bal67, Had73, VL66]. Input-Output [ABL75, Kre75, SP71b, Vid72, Wil71a, Bal67]. inputs [Kus68]. Instability [Bic68, Nol75, Sko74, SB74, Wil69]. Integer [Glo69]. Integrable [Ole74]. Integral [Hua72, RBR74, War70b, Cow73, Mae73, NW70, Vin69a, Vin69b, Vin69c, Vin70, Vin73]. Integrals [Bal74, Bal75, Ces75, Jac69a, Str66]. integrand [AD66]. Intensities [Fle71]. Interaction [Sch75a]. Interconnected [Bai65, Mic74]. Interior [Sas74]. Internal [WP74]. Interplanetary [Bre65]. Introduction [McM62]. Invariance [Onu71, Pen72, Ego66]. Invariant [FJ68, MS73, Pop72, SG63, Vid72, Sze62]. Invariants [WE75, Mor73]. Inverse [JK73]. Invertibility [De 74, Wil69]. investigation [Bla66]. Involving [Ban69, Wan75, Cow73, Goh66a]. Irrational [FG71]. Iterative [Mer64, PN70, Ros66, FY67, Gil66, HK66]. Ito [Bal74, Bal75, RBR74]. IV [Pel72]. Izv [Vin73].

J [CLS73, Luk73]. J. [Gil67]. Jeux [Ben74]. John [Geh74]. Joining [MP71a]. Joint [Gin67b]. Jump [BVW75a, BVW75b, Ris75, Sto73]. Junction [Mau75]. Justification [BF71, Bol66].

Kantorovich [GT73]. Keep [AS65]. kernel [Par62]. Kronecker [WE75]. Krotov [Gir69]. Kuhn [Gui69, Sch73, Zlo70].

Lag [Hal68b, Kra64a, Kra64b, Hal71]. Lagrange [CLS73, Ces65b, CLS71, Fal69, Nor71, Roc74]. Lagrangian [Nor67, Nor68]. Lagrangians [Man75]. lags [Ban68]. Laguerre [Doo64]. Large [Sch75a, Var66, BD68]. Large-Scale [Var66]. LaSalle [Hal64]. Laws [Ben71, Ris65]. learning [KD66]. Least [AS65, Kai72, NS75]. Least-Squares [NS75]. lemma [Jac67a]. Liapunov [Bur66, Kus67a, Lef65, Leh67, Yor71]. Lie [HH70]. Lienard [Bur65]. Life [DP72]. Life-Line [DP72]. Like [Ste74]. limit [Kul66, Kul67]. Limited [Bal74, Bal75, SR64]. Line [DP72]. linéaires [Ben74, Lob70, Lob76, Fau72]. Linear [AM69, Bab64, Bal72a, BJL75, BV75, Bha64, Bic68, Bra72, CR68, Chu74, CL66, Dan66, Dau71, DMM75, Doo64, EKM73, For75, Geh74, Gin67a, Gol66, GSG67, Hau74, HT70, Hey72, IMK72, IMK75, JK73, Kal63, KB70, Kus70, Kwa75, Lin73, Luk71, Luk72a, MN72, MF71, Nor71, NS75, O’M72, OK75, Pon74, Pop72, Psh66, Ris71b, Rob75, Roo69, Sak70, San75, Sca74b, SP71b, Smi64, SW75, SV64, TV75, Ven71, Vid72, WM70, WP74, Bal67, BJ70, Bha66, tC68, Con67, DIR73, Dau73, Fat68b, FJ68, Hau73a, HO73, HK66, Jac67b, Joh73, Kai69, Kus68, Med67, Mor73, Prè66, Rus66, Rus73, Sak66, SY71, Sap73, SS71, SM67, SA68, Sle74, Var70, Wah73, Wei68, Wil71b, Won67, Won73]. Linear-Quadratic [Sca74b]. Linearization [Kre73, Hol68, RM73]. Linearly [HH63, Lo75]. Lion [Fly73, Fly74a]. Local [Her74, DK70, Lor73b]. Locales [Lor73a, Lor73b]. locally [FFZ69]. Locus [Kra64a]. Logarithmic [FZ68]. Loop [Bru74, Hol74, Hol75, Kre69, Ste73, BD68]. Lower [Bau72, Ber74, RM73, Bru68, Ces71]. Lower-Dimensional [Bau72]. Lucien [Ano74]. Lur’e [Mey65]. Lyapunov [Bai65, Bel62, Kno75, Mey65, SB74, Wil71a, Zak69].

Man [Fly73, Fly74a]. Maneuver [Ho66]. Manifolds [Bro72, Lob74]. Many [Cas69]. mapping [OJ68]. Mappings [Dat70, RM73]. Maps [Kre75]. Markov [Kul66, Kul67, Kus64, Mil68, MS75, Pli75, Ris75, Sto73]. Markovian [Aka75, Chi69, Kak72]. Martingale [Bal72a]. Martingales [BVW75a, BVW75b, Won71]. Matching [Wol72]. Matematika [Vin73]. Mathematical [Kal63, MS71, Oka64, Sch67, Dan66, Gui69, Zlo70]. Matrices [And67, WE75, AM68, Paz67]. Matrix [ADT75, Hew75, Won68a, PD70, Won69].

Max [Fly74b, CM70a]. Max-Min [Fly74b, CM70a]. Maximum [IS75, Pli75, Wie72, Yor72, Cul66a, Hal66, HH66, KR66]. maximum-problems [KR66]. Mayer [Ces68]. Means [Doo64, LS75, Lin74b]. Measurability [Dat70, KO75]. measure [DS71]. measurement [Wil71b]. Measures [Kno75, Ris65, Sch65]. Mechanics [Sch75a, Mor66]. Memoriam [Ano74]. Method [AS65, Bai65, BF71, JM70, Kai72, Kün66, MT72, McC75, Mey73, Oka64, Sas72, Sas74, TAG70, Bal63, BM73, Bol66, Dan73, Gil67, GR69, Joh73, KP73, LH69, Leh67, PP73, Rub66]. Méthodes [Lor73a, Lor73b, MA74]. Methods [Ber75, Mey70b, Lor73b, MA74]. Mild [MT72]. Min [Fly74b, CM70a]. Minima [War67b]. Minimal [ABL75, Cul64, For75, PD70]. Minimal-Energy [Cul64]. Minimax [Sas74, Swo64, War65a, War70b]. minimisation [MA74]. Minimization [Ber75, Blu65, CM70b, DR68, PN70, TAG70, CCP66, DR67, MA74, Por68, Wah73]. Minimizing [Gol66]. Minimum [BT64, Joh73, Neu62, Neu65, Ris75, Gil66, Moy73, Rub66, Sno64]. Minimum-energy [Joh73]. Minimum-Fuel [Neu65]. minimum-norm [Rub66]. Missile [EH64]. Mixed [MN74, PH75]. Model [Wol72]. Modeling [Cul64, Roo75a, Roo75b]. Models [MP71b]. Modes [Har64]. Moment [Neu64]. Moments [Hau74, Zak67]. Monotone [Fre72, NN66, ZF68]. Motion [CH75, MS75]. multi [Fil67]. multi-valued [Fil67]. Multidimensional [Bau72, Dol73]. Multiperson [Pli73]. Multiple [Chy67, Goh66a, Hau73b]. Multiplicative [Hau74]. Multiplier [Nor71, Pon68, Roc74, WS74a]. Multipliers [Fal69, Fre68, RT73]. Multiserver [Hof75]. multistage [Wil71b]. multivalued [Jac69a, RM73]. Multivariable [For75, Smi64, Vid72, Vid75, WF69, WM70, WP74, Gil63, Gil69, Mor73, Won73].

Nash [Ben74, PR75, Pau75]. Necessary [ABI71, AT75, Aub70, Ban68, Bar71, BC72, Ber67, Ego67, GK72, Geh74, Goh66a, Hau73b, Kus72, MP71a, MP69, Ris70, Rus75, SRP69, SP71a, Sto73, Sur73, Vır70, Zol72, Jac69b]. Necessity [WS74a, Bru68]. Needed [Sus75]. Network [ADT75]. Networks [Dol74, DZ75, Ben66]. Neustadt [Ano74]. Neutral [BK72, Dat74]. Newton [McG65]. no [Vin73]. Noise [Fle71, Hau71, Hau74, Hol74, Zak69, Hau73a, Kus68, PD70, Won67]. Non [Fau72, Roo69, Lob70, Lob76]. Non-Gaussian [Roo69]. Nonautonomous [SH71]. Nonconvex [BT75, Fal69, Roc74]. Noncooperative [SP71a, SRP69]. nondifferentiable [BM73]. Nondiscounted [Kak72]. Nongradient [CM70b]. Nonlinear [ABI71, BSB63, Che70, Chu75, Cir69, GK74, HH70, HH63, Kre69, Kre73, Kre74, Kre75, Kün66, Lob74, Luk69, Luk72b, Luk74, Man63, Man75, Mar65, McC75, Neu64, Nol75, Roĭ70, Ros64, Sch65, Sch75b, SB74, WW69, Won64, Yor72, Zou66, Aro73, DK70, DeS73, Had73, Hal66, Hol68, KR66, Kro68, Lob70, Lob76, Luk73, Man66, NC68, Ros66, Rus69b, TV67, VL66, WD69]. Nonlinearities [Fre72, ZF68]. nonlinearity [Mae73, NN66]. Nonnegativity [Jac70, Mol75]. Nonsingular [Jac70, MP71a]. nonstandard [NY68]. Nonsymmetric [TV75]. nonzero [Var70]. Norm [Fre72, Rub66]. Normality [Haa75, MN72]. Normed [DR68, Gol66, Nor71]. Normed-Linear [Gol66]. Norms [PH75]. Note [EF75, FFA69, Kai72, SP71a, Sas72, Tro71, Gil67, Sle74]. nuclear [NY68]. Number [Sus75, Leh67]. Numbers [Sch75a]. Numerical [De 70, RBR74, BM73, Zou66].

Observability [Bra74, DM72a, DM72b, Jur70, Roĭ70, Sak75, Tri75, FJ68, Gil63, SM67, SA68]. Observable [AT74, Fle68, DV73]. One [LW75, Joh73, Mor66]. one-sided [Mor66]. Open [Hol74, Hol75, Ste73]. Operational [Fat64]. Operations [ADT75]. Operator [CL75, Bal63, Rus69b]. operator-valued [Rus69b]. Operatorial [Vır70]. Operators [Dav72a, FF69, Tri75]. Optima [SD71]. Optimal [AT74, Arm74, Bab64, Bal65, Bal68, Bar72a, Bar72b, Bar73, Ben71, BV75, Bis74, Bre65, Bri63, Bri64, Bru74, BBS69, BEL68, Ces65a, CS74, Cha64, Cha66, CL66, Chy67, Cul66b, Cul72, DM72b, Dem68, Dre64, Ego66, EH64, Fat64, Fle68, Gal69, Gin67a, Gin67b, Gol65a, GSG67, Gur64, Háj71, Hal68b, Hal71, Hau71, Her72, Hes65, HB63, Ho66, Hua72, JK73, JP64, Kel64, Kom70, Kre69, Kus65a, Kus67b, Lee63, Lei65, Lin74a, Lin74b, Luk69, MN74, Mal69, MP71a, Mir69, Mis65, NW70, Nis67, Nor67, Oğu63b, Ogu64, Pel72, Pon68, Por66, Psh66, Rus66, San75, SP71a, Sas72, SR64, Sch65, Sch72, Sny75, Vin69a, Vin69b, Vin69c, Vin70, Vin73, Wan75, War74, Won64, Won67, Zab75, Zol72, AD66, Ben66, Ben70]. optimal [Ber67, Bla66, Bri67, Bri68, Ces65b, Ces68, Ces71, CP66, Cul66a, Cul69, Dan73, DS71, Dav73, Fil62, FY67, Gam63, Gam65, Gri67, Jac67b, KD66, Kir63, Kir67, KP73, Kom68, Kus68, LL66, Man66, Med67, NY68, Nor68, PP73, Por68, Ris70, Ros66, RFL69, Rus69b, SRP69, SS71, Sch67, Sno64, Sto73, VW66]. Optimal-Thrust [Gur64]. Optimality [AT75, Bar71, FG70, GK72, Geh74, Kus65b, MP69, Bol66, Ego67, Hac73, Hay66, Jac69b, Bru74]. Optimization [Aub70, BT75, BSB63, De 70, Hal74, Kle74, Kus72, Mey70b, MP71b, Neu64, Pal66, Vır70, WS74a, WS74b, BJ70, BM73, Cow73, Hau73b, Lor73b, Neu66, Neu67, Sak66, Sur73]. Optimum [McG65, Oka64, Roo69, Wit68, Prè66]. Order [BJ71, GK72, Mau75, MP69, Rus75, San64, SH71, Tsu70, Joh73]. Ordinary [Man63, Onu71]. Origin [MDM74]. Orthogonal [Med67]. Oscillatory [Lel70, SY71]. Output [ABL75, Kre75, Mor71, SP71b, Vid72, Wil71a, Bal67, Had73, VL66]. Overrelaxation [Cry71].

paper [NW70, Vin73]. Parabolic

[AT75, Gal69, Mis65, Sak74, Sak75, Wan75, Zol72, Kus68, MB73]. Parallel [CM70b]. Parallelepipedon [Fat75]. Parameter [BT75, Bar71, BV75, Kus70, Kus72, LCB72, Pli75, Sas72, Zab74, Ego66, Ego67, SS71]. parametric [GR69]. Part [Roo75b, Rus71b]. Partial [AT75, ABL75, Ris71a, Sak74, Sak75, SG63, CM70a, Sur73, Sze62]. Partially [AT74, Fle68, DV73]. Particle [CH75]. parts [Leh67]. Patterns [Sko72]. Pénalisation [Lor73a, Lor73b]. Penalty [Ber75, EG74, Mur74, Oka64, Rus64, Sas72, Sas74, How73, Lor73b]. Performance [Wie72, Wit70, AD66, Rus73, Sak66]. Periodic [Dav72b, Kwa75]. Periodically [Bic68]. Periodicity [Hew75]. Périodiques [Fau72]. Person [SP71a, Sca74b, SRP69, Var70]. Personnes [Ben74]. Perturbations [Cul66b, Dau71, Dau73]. Perturbed [Gor72b, O’M72, OK75, San75, Zak69, DM69]. Phase [Fre72, Hua72, MN74, MN72, Rus64, Sca74a, Cul66a, Fre68, RT73]. Piecewise [BJ71, Gri67]. Plane [CH75, EZ71]. planning [Mil68]. Plans [Pel72]. Plates [Kom70]. Player [Cas69]. Point [Haa75, Mer64, MDM74, Gil67, Rub66]. Points [Ben74, CLS71, EKM73, Nor67, PR75, Rub75, VL69, CLS73, KR66, Nor68]. Poisson [Rus69a]. Polak [KP72]. Pole [MF71, MW70, WM70]. policies [Dav73].

Polyhedron [MDM74]. Polynomial [BJ71, WE75, Leh67]. Pontryagin [BC72, Ces65b, Hal66, IS75, Ris65]. Ports [Zem74]. Positive [And67, Bra72, AM68, AD66, Leh67, SY71, Sap73]. Positivity [Sko74]. pour [MA74]. Practical [WW69]. precision [KP73]. Prediction [KH69, Wei77]. Prediction-Correction [KH69, Wei77]. Preface [Bha72]. prescribed [Fre68, RT73]. presence [PD70]. Prices [Pel72]. Primal [Ber75]. Primal-Dual [Ber75]. Prime [Hey72]. Principle [Hal64, IS75, Ris65, SV64, Wie72, Yor72, Cul66a, Fat68b, Hal66, HH66, OJ68]. Principles [Ris75]. Probability [Kus64, Zak69, Paz67, Prè66]. Problem [Bar75b, Bar73, BJ71, Bre65, Cha64, Cry71, DMM75, FW74, Gar63, Gar64, Geh74, Gin67b, Gui65a, Gui65b, HH63, Her64, JK73, Mau75, Mey65, Mis65, Mis74, Neu64, O’M72, OK75, Oğu63b, PH75, SP71b, Sus72, TV75, Bal67, Bar69, Bri67, Bri68, CM70a, FY67, Goh66b, Kai69, Kir63, Por68, RFL69, Rub66, Sak66]. Problems [AT75, Aub70, Bal65, Ban69, BT75, Bar75a, Bar72a, Bar72b, Ber74, BBS69, Ces65a, CLS71, Cul66b, Cul72, Dan71, Dat73, De 70, Dem68, FG70, Gir69, Hal74, Hol75, Hua72, Jac70, Kel64, Kre74, Kus72, MN74, McS67, Mer64, Mey70a, Mie65, MN72, MH68, Oka64, Pet72, Psh66, Roc72, Rus75, Rus69a, Sak75, Sas72, Sas74, Sch72, Sch75c, TAG70, Vır70, War65a, War65b, War70a, War70b, Zol72, ARS67, Bal63, Ban68, Ben66, Ben70, Ber67, BM73, Bru68, CCP66, CE66, Ces65b, CLS73, CS70, Cow73, CP66, Cul66a, Cul69, Dan73, DM69, Ego66, Gam65, Gui69, Hau73b, Jac67b, Jac69b, KR66, Kro68, MB73, Moy73, NY68, Neu66, Neu67, Onu71, Par62, PP73, Ros66, SS71, Sno64, Sur73, Zlo70]. Procédé [Lor73a, Lor73b]. Procedure [Bar72b, PN70, Bar69, FY67, Gil66, KD66, Lor73b, MB73]. procedures [HK66]. Process [GSG67, IT75, Kak72, Lin74b, Rus69a, KD66]. Processes [Aka75, Bal74, Bal75, BVW75a, BVW75b, Cha66, Chi69, Chu75, IT75, Kus64, Kwa75, Lei65, Lin74a, Lin74b, LW75, Mar65, Ogu64, SL70, Bla66, Dan66, Ego66, Kir67, Kus67b, Mil68, NW70, Sto73, Vin69a, Vin69b, Vin69c, Vin70, Vin73]. product [MA74]. produit [MA74]. Program [SD71, Prè66]. Programming [ABI71, Ben66, Cry71, Gir69, Glo69, Gol65a, Kün66, MS71, Man75, McC75, Neu64, Ris71a, Ris75, Rob75, Roc74, VW66, Bar69, Bel66, Bol66, CE66, Dan66, DV73, Gui69, Kro68, Mor66, Ris70, TV67, Zlo70, Zou66]. Programs [Abr75, EH64, Fal69, GT73, WW69, GR69]. Projection [MT72, Med67]. Projections [Abr75]. Proof [BC72, Yor71]. Properties [Dat70, DZ75, Jur72, Onu71]. Property [CS74]. Provisions [Kle74]. Pseudo [Man65]. Pseudo-Convex [Man65]. Pseudodelays [War74]. publications [BEL68]. Pursuit [EF75, Fly74b, Mis65, Ogu64, Pet72, Por67, Rub72, Sak70]. Pursuit-Evasion [EF75, Por67, Sak70]. Pushdown [GS66].

Quadratic [Cry71, Dat73, Dat74, DMM75, KB70, LR69, Luk71, Luk72a, Mol75, Mor66, PN70, Rus73, Sca74b, TV75, Won71, Bar69, CE66, Gil66, GR69, Sak66, Wah73]. Quadratiques [Ben74]. Qualitative [DZ75]. Quasi [San75]. Quasi-Linear [San75]. Questions [AY70, Fil62].

R [NW70, Vin73]. Random

[IT75, Sch75a, Con67, OJ68, Prè66]. Randomly [AJ71]. Range [Ces75]. Raphson [McG65]. Rate [CC68, SR64]. Rates [SL70]. Rational [For75]. Reachability [DM69]. reachable [HO73]. reactor [NY68]. Real [And67, AM68, Leh67]. Realization [BB75, DIR74, Wei72, DeS73]. Realizations [ABL75, Kre75]. Realizing [Sko72]. Recoverability [SL70]. Recursive [Bal72a, Ris71b]. Region [GR75, For67]. regular [Hac73]. Regularity [MN72, WW69]. Regulation [CR68, Luk69, SW75, WP74, Rus66, Won73]. Regulator [BJ71, Cha64, O’M72, OK75, FY67]. Related [Sak75]. Relations [Gir69, Bal67, CP66, Kul66, Kul67, Mol73a]. Relaxed [McS67, Sch75c, War67a]. Relaxive [Zem74]. Relay [Oğu63a]. Remark [Chu65, SG63, Fat68b]. Remarks [Fat66, Jac67a, Tsu70, Zab74]. Representation [Aka75, BVW75a, Pli75, Won71, Str66]. Representations [Roo75a]. reproducing [Par62]. Resolution [De 74]. resolvent [Kai69]. Response [Har64]. Restricted [Hua72, Sca74a, Sch72, War67b, ZF68]. Restrictions [War70a]. Results [AM69, BVW75a, Buc66, CL75, Fly74b, WS74a]. Retardation [Oğu63a]. Retarded [BK72, Ogu64, LH69]. Reward [Pli75]. Ribière [KP72]. Riccati [Won69, Buc75, Buc77, CL75, Cur75b, Hew75, Mol73a, Mol73b, TV75, Won68a]. right [Fil67]. right-hand [Fil67]. Ritz [Dan73, MB73]. rockets [Ris66]. Root [Kra64a]. Root-Locus [Kra64a]. Rotational [LW75]. routing [Ben66]. Rubio [Gil67]. Rule [Nor71, WS74a].

Saddle [EKM73, Nor67, PR75, VL69, Nor68]. Salesman [Mur69]. Sampled [AJ71]. Satisfied [Kus64]. Scale [Var66]. Scheduling [Hof75]. search [Kro68]. Searching [Pau75]. Second [Bai65, BSB63, Jac70, MP69, Rus75, San64, Tsu70, Goh66b]. semi [Sto73]. semi-Markov [Sto73]. semicontinuity [Ces71, RM73]. Semiconvex [Wie72]. semidefinite [GR69]. Semigroup [Pli75]. semilinear [Sch67].

Sensitivity [Kre69, Por68]. Separable [WS74a, LL66]. Separation [Won68b]. Sequences [LwL67]. Sequential [ABL75, Chi69]. Sequentially [Rom69]. Set [Art75, CLS71, Dat70, PN70, Ste74, CLS73, Cul66a, DR67, Gil66, Paz67]. Set-Valued [Art75, Dat70]. Sets [Aub70, BK72, Ole74, HO73, Her67, Her68]. Several [BT64]. SIAM [CLS73, Luk73]. side [Fil67]. sided [Mor66]. simple [How73]. Simultanément [Lor73a, Lor73b]. simultaneously [Lor73b]. single [BD68, NN66]. single-loop [BD68]. Singular [Bre65, Haa75, Her64, Jac70, Kel64, Mau75, MP71a, Sno64, Goh66a, Goh66b, Hay66, Jac69b, Moy73]. Singularly [O’M72, OK75, San75]. Slender [CLS71, CLS73]. slope [ZF68]. slope-restricted [ZF68]. Small [Fle71, Hol74]. smooth [DR67]. smoothing [Kai69, Med67]. Solution [Bar73, Cry71, De 70, Dem68, Gar63, Gar64, Mer64, PH75, RBR74, Sak70, Bal63, CP66, Mol73b, Ros66, Sak66]. Solutions [DV71, DV75, Fat64, Fau72, Lel70, Ris71a, San64, Var67, Ces65b, Fil67, Rus69b, SS71, Zak67]. Solvability [FW74]. Solving [Bar72a, Bar72b, Oka64, FY67]. Some [Bur66, CL75, DK70, Dem68, DZ75, Dre64, Fat66, Fly74b, FG70, GT73, Jac67b, Lei65, Oka64, Pon65, Ros64, WW69, Wil71b, Won64, de 68, Gam65, Jac67a, TV67]. Soviet [BEL68]. Space [ABI71, Ack75, Bal72b, BK72, Coo75, Dat70, FG70, Gui69, Gui65a, Kak72, Mey70a, Neu65, Por67, Rob75, Sch72, Smi64, Ste74, TAG70, Tri75, Zab75, Zmo74, Cul66a, Fat68b, FFZ69, MA74, Sle74, Zlo70]. Space-Like [Ste74]. Spaces [Bal65, Bar75a, Bar75b, Bro72, De 74, DR68, For75, Gol66, Nor71, Tsu70, Vin74, CCP66, KR66, Par62]. spatial [Joh73]. specified [Ben70]. Spectral [Dav72b]. Spherical [GR75]. Spline [Dan71]. Splines [MS71]. Squares [AS65, Kai72, NS75]. Stability [Ack75, AM69, Bha64, Buc75, Buc77, Coo75, Dav72a, Dav72b, DMM75, Fre72, Gor72a, Gor72b, Hau73a, Kra64b, LaS62, Lef65, Mae73, Man63, Mic74, Mil72, NN66, NC68, Pen72, Ros64, Rox65, SH71, SG63, Ven71, Vid72, Vid75, Wei72, Wil69, ZF68, BD68, Bha66, tC68, FZ68, Fre68, FFZ69, Had73, Has68, LH69, Onu71, RT73, SA68, Sze62, VL66, WD69]. stabilizability [Sle74]. Stabilization [IMK72, WP74]. stabilizing [Mol73b]. Stable [Wil71a]. stage [KD66]. State [Bal72a, BJ71, Cul72, FG70, Hau71, Kak72, McG65, O’M72, OK75, Roc72, Roo69, Rus75, Sch72, Sch75b, Sch75c, Sny75, TV75, Wol72, Bal67, DeS73, Gil69, Joh73, Mil68, Won67]. State-Constrained [Cul72]. state-dependent [Won67]. state-relations [Bal67]. States [AKR67, BJL75, Bal67]. Stationary [Hau71, Hau74, Lin74a, Lin74b, Sko72, Sny75, Zak69, KR66, Kul66, Won67]. Statistical [Sch75a, Hol68]. Steepest [Gol65b, Vac66, Bal63]. steering [Aro73]. Stochastic [Aka75, Bal72b, Ben71, BV75, Bis74, Dre64, DV71, DV75, Fle71, Hof75, KO75, Kus65a, Kus65b, Kus72, Kwa75, Lin73, LW75, Roo69, SW75, Swo64, Wit68, Won64, Won68a, Won68b, Zab75, Ben70, CS70, DV73, Dav73, Kus67a, Kus67b, Paz67, Ris70, Str66, VW66, Won69, Zak67]. Strategies [SP71a, Ben70, Ber67, SRP69]. Strategy [Ho66, Ogu64]. Strict [De 74]. String [Mal69]. strong [Moy73, Rus69b]. Structural [Buc75, Buc77, Mor73]. Structure [DIR74, Hey72, MS73, SP71b, WF69, AM68, Wei68]. Structures [Wit71]. Subarcs [Kel64, MP71a]. Subject [CH75, Wah73]. Suboptimal [LwL67]. suboptimization [Bel66]. Subspaces [WE75, FJ68]. Sufficiency [Arm74]. Sufficient [Bar71, Bol66, FG70, Geh74, Jac70, Kus65b, Lee63, Man66, Moy73, DK70, Hac73, Ris70, Sto73]. sum [Var70]. Sure [AJ71]. Surveillance [ARS67, Rus69a]. survey [BEL68, Zou66]. Survival [Ste74]. Switching [Oğu63a]. symmetric [Rus66, Rus73]. Symmetries [Chu74]. Symmetry [IS75]. Synthesis [Bab64, MP71b, Mir69, Mor71, Bri67, Bri68, Hac73]. System [And67, AM69, AY70, Bro72, DV71, DV75, Gin67a, Hua72, Mor71, OJ68, Bel66, Kus68, Por68, Won67]. Systematic [Cry71]. Systèmes [Fau72, Lob70, Lob76]. Systems [Ack75, AJ71, AT75, AM75, Bab64, Bai65, BJL75, Bar71, Bau72, BV75, Bha64, Bic68, Bra72, BSB63, CR68, Che70, Chu74, CL66, Chy67, Cir69, Cul64, Cur75b, DIR74, Dau71, Dav72b, De 74, DM72a, DM72b, DMM75, DC72, Dol73, Doo64, Dre64, Fat68a, For75, FG71, Fre72, Gal69, Hal68b, Hau74, HT70, Hey72, Hof75, IMK72, IMK75, JP64, Jur70, Jur72, Kal63, Kra64b, Kre69, Kre73, Kus70, LCB72, Lin73, Lo75, Lob74, Luk69, LR69, Luk72b, Mic74, MF71, MS75, MS73, Neu62, Nol75, NS75, Oğu63a, Oğu63b, Ogu64, Pen72, Pop72, Por66, RM68, Ris65, Ris75, Ris71b, Roĭ70, Roo69, Roo75a, Roo75b, Rox65, San64, San75, Sch75a, Sch75b, SP71b, Sko72, Sko74, SW75, SV64, SB74, Sus72, Swo64, SG63, Var66, Ven71]. Systems [Vid72, Vid75, Wan75, Wei67, Wei72, Wil71a, Wit70, WF69, WM70, WP74, Zab74, Zab75, Zak69, Zmo74, Aro73, BD68, Bal67, Bha66, BEL68, tC68, Con67, DIR73, Dau73, DV73, DK70, DeS73, DM69, Ego66, Ego67, Fat68b, FJ68, For67, FZ68, Fre68, Gil63, Gil69, Had73, Hal71, Hal66, Hau73a, HO73, HH66, Hol68, Joh73, LH69, LL66, Lob70, Lob76, Luk73, Man66, Mor73, NC68, RFL69, Rus66, Rus73, Sak66, SY71, Sap73, Sch67, SM67, SA68, Sle74, Sze62, VL66, Wei68, Won73, WD69, ZF68].

Target [BK72]. Technique [Bal68, Pea66a]. telephone [Ben66]. Temperature [Fat75]. Tenability [Sch75b]. Terminal [BT64, Ho66, Ste74, Joh73]. Their [Pet72].

Theorem [AT74, Ces65a, CLS71, Ces74, EG74, Kre74, Nis67, Vin74, Won68b, Yor71, CLS73, Cow73, HO73, Hay66]. Theorems [Bau72, Ber74, Bur66, Gor72a, KO75, de 68, Ces65b, Ces68, Ces71, Has68, Jac67b, Kus67a]. theoretic [Bal63, Kul66, Kul67]. Theory [And67, Arb65, Arm74, BB75, Bar73, BB64, Bro72, Buc66, Cas69, CS74, Cha66, Cul64, DIR74, Dav72b, EG74, FF69, Gor72b, Háj71, HH70, Hes65, IMK75, IT75, Kle74, Lee63, Mir69, Neu65, Oka64, Pli75, Rus71b, Sch65, Smi64, DIR73, Ego66, Fil62, FFZ69, Gam65, Kir67, Neu66, Neu67, Onu71, Sch67, Wei68]. Thin [Kom70]. Thrust [Gur64, Ris66]. Tight [CC68]. Time [Ack75, AM75, Bic68, Bru74, Chi69, Chu75, CL66, DMM75, Fat64, GSG67, Háj71, Hal68b, Har64, Her72, IMK75, JP64, Kak72, Kra64a, Kra64b, LCB72, Lin74a, Lin74b, Mey70a, MF71, Nol75, Oğu63b, Pop72, Roo75b, Ros64, SR64, Sko72, Swo64, SG63, Ven71, Vid72, Wan75, Wei72, Zab74, Zab75, Bal67, Ban68, tC68, DK70, FZ68, Fre68, FY67, Had73, Hal71, Mil68, NC68, Ris70, SM67, Sze62, VL66, WD69]. Time-Dependent [Ros64]. Time-Invariant [Pop72, Vid72]. Time-Optimal [Bru74, Fat64, Háj71, Her72, FY67]. time-variable [SM67]. Time-Varying [Ack75, AM75, Bic68, IMK75, Nol75, Roo75b, Sko72, Bal67, tC68, DK70, FZ68, Fre68, Had73, NC68, VL66, WD69]. Topological [Pet72]. topology [Rus69b]. totally [Hay66]. Tours [Mur69]. Tracer [Pap73]. Tracking [FG71, Won73]. Trajectories [Gur64, Neu65, BJ70]. Trajectory [Kel64]. Transfer [FG71, HT70]. Transform [FF69]. Transformation [Kel64]. transverse [Kom68]. Traveling [Mur69]. Tucker [Gui69, Sch73, Zlo70]. Turn [GS66]. Two [LS75, Mer64, Sch67, ARS67].

Two-dimensional [ARS67]. Two-Point [Mer64]. Type [BK72, Oğu63a, Sak74, Sak75, Ces68, Hal66]. Types [Dre64].

Ucebn [Vin73]. ultimate [Zak67]. Unattained [Abr75]. Unbounded [CLS71, War65b, CLS73]. Uncertain [Pen72, Wit70]. uncertainty [VW66, Wil71b]. Unconstrained [CM70b, Dat73, Man75]. Unilateral [War65a, War70b]. uniqueness [HO73]. Upper [CC68]. Use [Kle74, Wol72]. Used [NS75]. Using [Bal74, Bal75, BSB63, SB74, Cry71, Joh73, Lor73b, SY71]. Utilisant [Lor73a, Lor73b].

V [Vin73]. V. [NW70]. Validity [Mey70b]. Value [GR75, Mer64, Rus71a, Rus71b, Ste72, RT73]. Valued [Art75, Dat70, Fil67, Rus69b]. Values [SD71]. Variable [MN74, San64, Ban68, Joh73, SM67]. Variables [Aka75, Roo69, Goh66a, Sch67, Vac66]. Variation [BSB63, Jac70, Ven71, Won71, FZ68, Goh66b, Lor73b]. Variational [Ban69, Gui65b, Hes65, McS67, War65b, Neu66, Neu67]. Variations [Lor73a, War65a, Lor73b]. Varying [Ack75, AM75, Bic68, IMK75, Nol75, Roo75b, Sko72, Ven71, Bal67, tC68, DK70, FZ68, Fre68, Had73, NC68, VL66, WD69]. Vector [Bel62, For75, Kno75, LS75, Pli75, Sch65, Hay66]. vectors [Paz67]. Vehicle [Smi64]. Via [FG71, BJ71, HH70, MS71, WE75]. Vibrating [Kom70, Mal69]. vibration [Kom68]. Vinokurov [NW70, Vin73]. Volterra [Mae73]. Vyss [Vin73].

W [Ano74]. Wave [GR75, Rus71a, Rus71b]. Weak [Art75, Ole74, Ris71a]. Weakened [Gui65b]. Weber [PH75]. Weighting [Sko72]. Which [Aub70, Cul66a]. White [IT75, Kus68]. whose [Paz67, Ris65]. without [Ces74, Hal74]. Wolfe [CC68, Mey74].

Zaved [Vin73]. zeros [Leh67].

References

Abrams:1971:NPC

[ABI71] Robert A. Abrams and Adi Ben-Israel. Nonlinear programming in complex space: Necessary conditions. SIAM Journal on Control, 9 (4):606–620, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Anderson:1975:SCM

[ABL75] B. M. Anderson, F. M. Brasch, Jr., and P. V. Lopresti. Sequential construction of minimal partial realizations from finite input-output data. SIAM Journal on Control, 13(3): 552–571, May 1975. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Abrams:1975:PCP

[Abr75] Robert A. Abrams. Projections of convex programs with unattained infima. SIAM Journal on Control, 13(3):706–718, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Acker:1975:SCT

[Ack75] Andrew Acker. Stability criteria for time-varying systems in Hilbert space. SIAM Journal on Control, 13 (6):1156–1171, November 1975. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Anvari:1966:EOC

[AD66] M. Anvari and R. F. Datko. The existence of optimal controls for a performance index with a positive integrand. SIAM Journal on Control, 4(2):372–381, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Anderson:1975:MOI

[ADT75] William N. Anderson, Jr., Richard J. Duffin, and George E. Trapp. Matrix operations induced by network connections. SIAM Journal on Control, 13(2):446–461, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Agniel:1971:ASB

[AJ71] R. G. Agniel and E. I. Jury. Almost sure boundedness of randomly sampled systems. SIAM Journal on Control, 9(3):372–384, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Akaike:1975:MRS

[Aka75] Hirotugu Akaike. Markovian representation of stochastic processes by canonical variables. SIAM Journal on Control, 13(1):162–173, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Akilov:1967:ESE

[AKR67] G. P. Akilov, L. V. Kantorovich, and G. Sh. Rubinshtein. Extremal states and extremal controls. SIAM Journal on Control, 5(4):600–608, ????

1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Anderson:1968:ASG

[AM68] B. D. O. Anderson and J. B. Moore.

Algebraic structure of generalized positive real matrices. SIAM Journal on Control, 6(4):615–624, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Anderson:1969:NRL

[AM69] B. D. O. Anderson and J. B. Moore.

New results in linear system stability. SIAM Journal on Control, 7 (3):398–414, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Arbib:1975:TVS

[AM75] Michael A. Arbib and Ernest G.

Manes. Time-varying systems. SIAM Journal on Control, 13(6): 1252–1270, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Anderson:1967:STC

[And67] B. D. O. Anderson. A system theory criterion for positive real matrices. SIAM Journal on Control, 5 (2):171–182, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Anonymous:1974:LWN

[Ano74] Anonymous. Lucien W. Neustadt — in memoriam. SIAM Journal on Control, 12(2):179–183, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Arbib:1965:CFA

[Arb65] Michael A. Arbib. A common framework for automata theory and control theory. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2):206–222, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Armstrong:1974:EOC

[Arm74] Gerald M. Armstrong. An extension of an optimal control sufficiency theory. SIAM Journal on Control, 12(4):650–654, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Aronsson:1973:GCB

[Aro73] Gunnar Aronsson. Global controllability and bang-bang steering of certain nonlinear systems. SIAM Journal on Control, 11(4):607–619, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Antelman:1967:SPT

[ARS67] G. R. Antelman, C. B. Russell, and I. R. Savage. Surveillance problems: Two-dimensional with continuous surveillance. SIAM Journal on Control, 5(2):245–267, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Artstein:1975:WCS

[Art75] Zvi Artstein. Weak convergence of set-valued functions and control. SIAM Journal on Control, 13(4): 865–878, July 1975. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Albert:1965:MCL

[AS65] Arthur Albert and Robert W. Sittler. A method for computing least squares estimators that keep up with the data. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3):384–417, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Ahmed:1974:ETO

[AT74] N. U. Ahmed and K. L. Teo. An existence theorem on optimal control of partially observable diffusions. SIAM Journal on Control, 12 (3):351–355, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ahmed:1975:NCO

[AT75] N. U. Ahmed and K. L. Teo. Necessary conditions for optimality of Cauchy problems for parabolic partial differential systems. SIAM Journal on Control, 13(5):981–993, August 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Aubin:1970:CSC

[Aub70] Jean Pierre Aubin. Characterization of the sets of constraints for which the necessary conditions for optimization problems hold. SIAM Journal on Control, 8(2):148–162, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Aoki:1970:CCQ

[AY70] M. Aoki and P. C. Yue. On certain convergence questions in system identification. SIAM Journal on Control, 8(2):239–256, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Babunashvili:1964:SLO

[Bab64] T. G. Babunashvili. The synthesis of linear optimal systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):261–265, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Bailey:1965:ALS

[Bai65] F. N. Bailey. The application of Lyapunov’s second method to interconnected systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3): 443–462, ???? 1965. CODEN JSM-CAL. ISSN 0887-4603.

Balakrishnan:1963:OTF

[Bal63] A. V. Balakrishnan. An operator theoretic formulation of a class of control problems and a steepest descent method of solution. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(2):109–127, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Balakrishnan:1965:OCP

[Bal65] A. V. Balakrishnan. Optimal control problems in Banach spaces. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):152–180, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Balakrishnan:1967:PDS

[Bal67] A. V. Balakrishnan. On the problem of deducing states and staterelations from input-output relations for linear time-varying systems.

SIAM Journal on Control, 5(3):309– 325, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Balakrishnan:1968:NCT

[Bal68] A. V. Balakrishnan. On a new computing technique in optimal control. SIAM Journal on Control, 6(2):149– 173, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Balakrishnan:1972:MAL

[Bal72a] A. V. Balakrishnan. Martingale approach to linear recursive state estimation. SIAM Journal on Control, 10(4):754–766, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Balakrishnan:1972:SCF

[Bal72b] A. V. Balakrishnan. Stochastic control: a function space approach. SIAM Journal on Control, 10(2): 285–297, ???? 1972. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Balakrishnan:1974:AII

[Bal74] A. V. Balakrishnan. On the approximation of Ito integrals using bandlimited processes. SIAM Journal on Control, 12(2):237–251, May 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See errata [Bal75].

Balakrishnan:1975:EAI

[Bal75] A. V. Balakrishnan. Errata: On the approximation of Ito integrals using band-limited processes. SIAM Journal on Control, 13(4):975, ???? 1975.

CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Bal74].

Banks:1968:NCC

[Ban68] H. T. Banks. Necessary conditions for control problems with variable time lags. SIAM Journal on Control, 6(1):9–47, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Banks:1969:VPI

[Ban69] H. T. Banks. Variational problems involving functional differential equations. SIAM Journal on Control, 7(1):1–17, February 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Barr:1969:ECP

[Bar69] Robert O. Barr. An efficient computational procedure for a generalized quadratic programming problem. SIAM Journal on Control, 7 (3):415–429, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Barnes:1971:NSO

[Bar71] Earl R. Barnes. Necessary and sufficient optimality conditions for a class of distributed parameter control systems. SIAM Journal on Control, 9 (1):62–82, February 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Barnes:1972:GCA

[Bar72a] Earl R. Barnes. A geometrically convergent algorithm for solving optimal control problems. SIAM Journal on Control, 10(3):434–443, ????

1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Barnes:1972:PSC

[Bar72b] Earl R. Barnes. A procedure for solving certain dual optimal control problems. SIAM Journal on Control, 10(2):377–392, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Barnes:1973:SDP

[Bar73] Earl R. Barnes. Solution of a dual problem in optimal control theory. SIAM Journal on Control, 11(2): 344–357, ???? 1973. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Barbu:1975:CCP

[Bar75a] Viorel Barbu. Convex control problems of Bolza in Hilbert spaces. SIAM Journal on Control, 13(4): 754–771, July 1975. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Barbu:1975:CPB

[Bar75b] Viorel Barbu. On the control problem of Bolza in Hilbert spaces. SIAM Journal on Control, 13(5): 1062–1076, August 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Baum:1972:ETM

[Bau72] Richard F. Baum. Existence theorems for multidimensional control systems with lower-dimensional controls. SIAM Journal on Control, 10 (4):623–638, November 1972. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bellman:1964:ACT

[BB64] Richard Bellman and Richard Bucy. Asymptotic control theory. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1):11–18, ???? 1964. CO-DEN JSMCAL. ISSN 0887-4603.

Baras:1975:FID

[BB75] J. S. Baras and R. W. Brockett. H^{2}- functions and infinite-dimensional realization theory. SIAM Journal on Control, 13(1):221–241, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Budak:1969:DAO

[BBS69] B. M. Budak, E. M. Berkovich, and E. N. Solov′eva. Difference approximations in optimal control problems. SIAM Journal on Control, 7 (1):18–31, February 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Baum:1972:RPP

[BC72] R. F. Baum and L. Cesari. On a recent proof of Pontryagin’s necessary conditions. SIAM Journal on Control, 10(1):56–75, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Baker:1968:ASL

[BD68] R. A. Baker and C. A. Desoer. Asymptotic stability in the large of a class of single-loop feedback systems. SIAM Journal on Control, 6(1):1–8, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bellman:1962:VLF

[Bel62] Richard Bellman. Vector Lyapunov functions. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(1): 32–34, ???? 1962. CODEN JSM-CAL. ISSN 0887-4603.

Bellman:1966:DPS

[Bel66] Richard Bellman. Dynamic programming, system identification, and suboptimization. SIAM Journal on Control, 4(1):1–5, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Butkovsky:1968:OCD

[BEL68] A. G. Butkovsky, A. I. Egorov, and K. A. Lurie. Optimal control of distributed systems (a survey of Soviet publications). SIAM Journal on Control, 6(3):437–476, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Benes:1966:PCP

[Ben66] V. E. Beneš. Programming and control problems arising from optimal routing in telephone networks. SIAM Journal on Control, 4(1):6–18, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Benes:1970:EOS

[Ben70] V. E. Beneš. Existence of optimal strategies based on specified information, for a class of stochastic decision problems. SIAM Journal on Control, 8(2):179–188, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Benes:1971:EOS

[Ben71] V. E. Benes. Existence of optimal stochastic control laws. SIAM Journal on Control, 9(3):446–472, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bensoussan:1974:PNC

[Ben74] A. Bensoussan. Points de Nash dans le cas de fonctionnelles quadratiques et jeux differentiels linéaires à N personnes. SIAM Journal on Control, 12(3):460–499, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Berkovitz:1967:NCO

[Ber67] Leonard D. Berkovitz. Necessary conditions for optimal strategies in a class of differential games and control problems. SIAM Journal on Control, 5(1):1–24, ???? 1967. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Berkovitz:1974:ELC

[Ber74] Leonard D. Berkovitz. Existence and lower closure theorems for abstract control problems. SIAM Journal on Control, 12(1):27–42, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bertsekas:1975:CPD

[Ber75] Dimitri P. Bertsekas. Combined primal-dual and penalty methods for constrained minimization. SIAM Journal on Control, 13(3):521–544, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bergen:1971:JDF

[BF71] A. R. Bergen and R. L. Franks. Justification of the describing function method. SIAM Journal on Control, 9(4):568–589, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bhatia:1964:ESL

[Bha64] Nam P. Bhatia. On exponential stability of linear differential systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):181–191, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Bhatia:1966:AES

[Bha66] Nam P. Bhatia. Addendum: “On exponential stability of linear differential systems”. SIAM Journal on Control, 4(4):695–697, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bhatia:1972:P

[Bha72] Nam P. Bhatia. Preface. SIAM Journal on Control, 10(2):vii, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bickart:1968:IPT

[Bic68] Theodore A. Bickart. Instability of periodically time-varying linear control systems. SIAM Journal on Control, 6(1):48–57, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bismut:1974:EOS

[Bis74] Jean-Michel Bismut. An example of optimal stochastic control with constraints. SIAM Journal on Control, 12(3):401–418, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Banks:1970:OTL

[BJ70] H. T. Banks and Marc Q. Jacobs. The optimization of trajectories of linear functional differential equations. SIAM Journal on Control, 8 (4):461–488, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bosarge:1971:EBH

[BJ71] W. E. Bosarge, Jr. and O. G. Johnson. Error bounds of high order accuracy for the state regulator problem via piecewise polynomial approximations. SIAM Journal on Control, 9 (1):15–28, February 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Banks:1975:CCS

[BJL75] H. T. Banks, Marc Q. Jacobs, and C. E. Langenhop. Characterization of the controlled states in W_{2}^{(1)} of linear hereditary systems. SIAM Journal on Control, 13(3):611–649, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Banks:1972:CFD

[BK72] H. T. Banks and G. A. Kent. Control of functional differential equations of retarded and neutral type to target sets in function space. SIAM Journal on Control, 10(4):567–593, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Blaquiere:1966:FIG

[Bla66] A. Blaquière. Further investigation into the geometry of optimal processes. SIAM Journal on Control, 4(1):19–33, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Blum:1965:MFE

[Blu65] E. K. Blum. Minimization of functionals with equality constraints. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2):299–316, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Bertsekas:1973:DNM

[BM73] Dimitri P. Bertsekas and Sanjoy K.

Mitter. A descent numerical method for optimization problems with nondifferentiable cost functionals. SIAM Journal on Control, 11(4):637–652, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Boltyanskii:1966:SCO

[Bol66] V. G. Boltyanskii. Sufficient conditions for optimality and the justification of the dynamic programming method. SIAM Journal on Control, 4(2):326–361, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Brammer:1972:CLA

[Bra72] Robert F. Brammer. Controllability in linear autonomous systems with positive controllers. SIAM Journal on Control, 10(2):339–353, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Brammer:1974:GCO

[Bra74] Robert F. Brammer. Geometrically constrained observability. SIAM Journal on Control, 12(3):449–459, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Breakwell:1965:DSP

[Bre65] J. V. Breakwell. A doubly singular problem in optimal interplanetary guidance. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1): 71–77, ???? 1965. CODEN JSM-CAL. ISSN 0887-4603.

Bridgland:1963:EOF

[Bri63] T. F. Bridgland, Jr. On the existence of optimal feedback controls. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3):261–274, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Bridgland:1964:EOF

[Bri64] T. F. Bridgland, Jr. On the existence of optimal feedback controls. II. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):137–150, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Bridgland:1967:PAS

[Bri67] T. F. Bridgland, Jr. On the problem of approximate synthesis of optimal controls. SIAM Journal on Control, 5(3):326–344, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See errata [Bri68].

Bridgland:1968:EPA

[Bri68] T. F. Bridgland, Jr. Errata: “On the problem of approximate synthesis of optimal controls”. SIAM Journal on Control, 6(1):58, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Bri67].

Brockett:1972:STG

[Bro72] R. W. Brockett. System theory on group manifolds and coset spaces. SIAM Journal on Control, 10(2): 265–284, ???? 1972. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Brunovsky:1968:NCC

[Bru68] Pavol Brunovský. On the necessity of a certain convexity condition for lower closure of control problems. SIAM Journal on Control, 6(2):174– 185, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Brunovsky:1974:CLT

[Bru74] Pavol Brunovský. The closed-loop time-optimal control. I: Optimality. SIAM Journal on Control, 12(4): 624–634, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Breakwell:1963:OCN

[BSB63] John V. Breakwell, Jason L. Speyer, and Arthur E. Bryson. Optimization and control of nonlinear systems using the second variation. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(2):193–223, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Breakwell:1964:MEC

[BT64] J. V. Breakwell and F. Tung. Minimum effort control of several terminal components. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3): 295–316, ???? 1964. CODEN JSM-CAL. ISSN 0887-4603.

Baranger:1975:NOP

[BT75] J. Baranger and R. Temam. Nonconvex optimization problems depending on a parameter. SIAM Journal on Control, 13(1):146–152, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bucy:1966:NRA

[Buc66] R. S. Bucy. New results in asymptotic control theory. SIAM Journal on Control, 4(3):397–402, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bucy:1975:SSR

[Buc75] R. S. Bucy. Structural stability for the Riccati equation. SIAM Journal on Control, 13(4):749–753, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Buc77].

Bucy:1977:ESS

[Buc77] R. S. Bucy. Erratum: Structural stability for the Riccati equation. SIAM Journal on Control and Optimization, 15(5):904, August 1977. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Buc75].

Burton:1965:GLE

[Bur65] T. A. Burton. The generalized Lienard equation. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2): 223–230, ???? 1965. CODEN JSM-CAL. ISSN 0887-4603.

Burton:1966:SLT

[Bur66] T. A. Burton. Some Liapunov theorems. SIAM Journal on Control, 4(3):460–465, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Bensoussan:1975:OCS

[BV75] A. Bensoussan and M. Viot. Optimal control of stochastic linear distributed parameter systems. SIAM Journal on Control, 13(4):904–926, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Boel:1975:MJPa

[BVW75a] R. Boel, P. Varaiya, and E. Wong.

Martingales on jump processes. i: Representation results. SIAM Journal on Control, 13(5):999–1021, August 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Boel:1975:MJPb

[BVW75b] R. Boel, P. Varaiya, and E. Wong.

Martingales on jump processes. II: Applications. SIAM Journal on Control, 13(5):1022–1061, August 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Case:1969:TTM

[Cas69] James H. Case. Toward a theory of many player differential games. SIAM Journal on Control, 7(2):179– 197, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Canon:1968:TUB

[CC68] M. D. Canon and C. D. Cullum. A tight upper bound on the rate of convergence of Frank–Wolfe algorithm. SIAM Journal on Control, 6(4):509– 516, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Canon:1966:CMP

[CCP66] M. Canon, C. Cullum, and E. Polak. Constrained minimization problems in finite-dimensional spaces. SIAM Journal on Control, 4(3):528–547, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Canon:1966:NAC

[CE66] M. D. Canon and James H. Eaton. A new algorithm for a class of quadratic programming problems with application to control. SIAM Journal on Control, 4(1):34–45, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cesari:1965:ETP

[Ces65a] Lamberto Cesari. An existence theorem in problems of optimal control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):7–22, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Cesari:1965:ETO

[Ces65b] Lamberto Cesari. Existence theorems for optimal solutions in Pontryagin and Lagrange problems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3):475–498, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Cesari:1968:ETO

[Ces68] Lamberto Cesari. Existence theorems for optimal controls of the Mayer type. SIAM Journal on Control, 6(4):517–552, ???? 1968. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cesari:1971:CLC

[Ces71] Lamberto Cesari. Closure, lower closure, and semicontinuity theorems in optimal control. SIAM Journal on Control, 9(2):287–315, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cesari:1974:ETC

[Ces74] Lamberto Cesari. An existence theorem without convexity conditions. SIAM Journal on Control, 12(2): 319–331, May 1974. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cesari:1975:CRC

[Ces75] Lamberto Cesari. Convexity of the range of certain integrals. SIAM Journal on Control, 13(3):666–676, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cockayne:1975:PMP

[CH75] E. J. Cockayne and G. W. C. Hall. Plane motion of a particle subject to curvature constraints. SIAM Journal on Control, 13(1):197–220, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chang:1964:ORP

[Cha64] Albert Chang. An optimal regulator problem. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):220–233, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Chang:1966:GTO

[Cha66] Sheldon S. L. Chang. General theory of optimal processes. SIAM Journal on Control, 4(1):46–55, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cheprasov:1970:CNS

[Che70] V. A. Cheprasov. On controllability of nonlinear systems. SIAM Journal on Control, 8(1):113–123, February 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chitgopekar:1969:CTM

[Chi69] S. S. Chitgopekar. Continuous time Markovian sequential control processes. SIAM Journal on Control, 7 (3):367–389, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chu:1965:RCC

[Chu65] Hsin Chu. A remark on complete controllability. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3):

439–442, ???? 1965. CODEN JSM-CAL. ISSN 0887-4603.

Chukwu:1974:SAL

[Chu74] Ethelbert Nwakuche Chukwu. Symmetries of autonomous linear control systems. SIAM Journal on Control, 12(3):436–448, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chukwu:1975:FTC

[Chu75] Ethelbert N. Chukwu. Finite time controllability of nonlinear control processes. SIAM Journal on Control, 13(4):807–816, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chyung:1967:OSM

[Chy67] Dong Hak Chyung. Optimal systems with multiple cost functionals. SIAM Journal on Control, 5(3):345– 351, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cirina:1969:BCN

[Cir69] Marco Cirinà. Boundary controllability of nonlinear hyperbolic systems. SIAM Journal on Control, 7(2):198–212, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chyung:1966:LOS

[CL66] D. H. Chyung and E. Bruce Lee. Linear optimal systems with time delays. SIAM Journal on Control, 4 (3):548–575, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Casti:1975:SNA

[CL75] J. Casti and L. Ljung. Some new analytic and computational results for operator Riccati equations. SIAM Journal on Control, 13(4):817–826, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cesari:1971:ETL

[CLS71] L. Cesari, J. R. La Palm, and D. A. Sanchez. An existence theorem for Lagrange problems with unbounded controls and a slender set of exceptional points. SIAM Journal on Control, 9(4):590–605, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [CLS73].

Cesari:1973:EET

[CLS73] L. Cesari, J. R. La Palm, and D. A. Sánchez. Erratum: “An existence theorem for Lagrange problems with unbounded controls and a slender set of exceptional points” (SIAM J. Control 9 (1971), 590–605). SIAM Journal on Control, 11(4):677, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [CLS71].

Cea:1970:EMM

[CM70a] Jean Céa and Kazimierz Malanowski. An example of a max-min problem in partial differential equations. SIAM Journal on Control, 8(3):305–316, August 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chazan:1970:NPA

[CM70b] D. Chazan and W. L. Miranker.

Nongradient and parallel algorithm for unconstrained minimization. SIAM Journal on Control, 8(2):207– 217, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Connors:1967:CDL

[Con67] Michael M. Connors. Controllability of discrete, linear, random dynamical systems. SIAM Journal on Control, 5(2):183–210, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cook:1975:CCS

[Coo75] P. A. Cook. Circle criteria for stability in Hilbert space. SIAM Journal on Control, 13(3):593–610, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cowles:1973:ETO

[Cow73] David E. Cowles. An existence theorem for optimization problems involving integral equations. SIAM Journal on Control, 11(4):595–606, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cullum:1966:ERC

[CP66] C. D. Cullum and E. Polak. Equivalence relations for the classification and solution of optimal control problems. SIAM Journal on Control, 4 (3):403–420, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chang:1968:RII

[CR68] A. Chang and J. Rissanen. Regulation of incompletely identified linear systems. SIAM Journal on Control, 6(3):327–348, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cryer:1971:SQP

[Cry71] Colin W. Cryer. The solution of a quadratic programming problem using systematic overrelaxation. SIAM Journal on Control, 9(3):385–392, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chen:1970:DPD

[CS70] I-Ngo Chen and C. L. Sheng. The decision problems of definite stochastic automata. SIAM Journal on Control, 8(1):124–134, February 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cesari:1974:CPO

[CS74] L. Cesari and M. B. Suryanarayana. Convexity and property (Q) in optimal control theory. SIAM Journal on Control, 12(4):705–720, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Culver:1964:ATM

[Cul64] W. J. Culver. An analytic theory of modeling for a class of minimalenergy control systems (disturbancefree case). Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):267–294, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Cullum:1966:MPO

[Cul66a] Jane Cullum. A maximum principle for optimal control problems in which the phase space constraint set is closed. SIAM Journal on Control, 4(3):488–504, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cullum:1966:POC

[Cul66b] Jane Cullum. Perturbations of optimal control problems. SIAM Journal on Control, 4(3):473–487, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cullum:1969:DAC

[Cul69] Jane Cullum. Discrete approximations to continuous optimal control problems. SIAM Journal on Control, 7(1):32–49, ???? 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Cullum:1972:FDA

[Cul72] Jane Cullum. Finite-dimensional approximations of state-constrained continuous optimal control problems. SIAM Journal on Control, 10 (4):649–670, November 1972. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Curtain:1975:IDF

[Cur75a] Ruth F. Curtain. Infinitedimensional filtering. SIAM Journal on Control, 13(1):89–104, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Curtain:1975:IDR

[Cur75b] Ruth F. Curtain. The infinite dimensional Riccati equation with applications to affine hereditary differential systems. SIAM Journal on Control, 13(6):1130–1143, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Dantzig:1966:LCP

[Dan66] George B. Dantzig. Linear control processes and mathematical programming. SIAM Journal on Control, 4(1):56–60, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Daniel:1971:CDC

[Dan71] James W. Daniel. Convergence of a discretization for constrained spline function problems. SIAM Journal on Control, 9(1):83–96 (or 83–86??), February 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Daniel:1973:RGM

[Dan73] James W. Daniel. The Ritz–Galerkin method for abstract optimal control problems. SIAM Journal on Control, 11(1):53–63, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Datko:1970:MPS

[Dat70] Richard Datko. Measurability properties of set-valued mappings in a Banach space. SIAM Journal on Control, 8(2):226–238, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Datko:1973:UCP

[Dat73] Richard Datko. Unconstrained control problems with quadratic cost. SIAM Journal on Control, 11(1):32– 52, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Datko:1974:NAF

[Dat74] R. Datko. Neutral autonomous functional equations with quadratic cost. SIAM Journal on Control, 12(1):70– 82, February 1974. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Dauer:1971:PLC

[Dau71] Jerald P. Dauer. Perturbations of linear control systems. SIAM Journal on Control, 9(3):393–400, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Dau73].

Dauer:1973:EPL

[Dau73] Jerald P. Dauer. Errata: “Perturbations of linear control systems”. SIAM Journal on Control, 11(2):395, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Dau71].

Davis:1972:FOE

[Dav72a] Jon H. Davis. Fredholm operators, encirclements, and stability criteria. SIAM Journal on Control, 10(4): 608–622, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Davis:1972:SCD

[Dav72b] Jon H. Davis. Stability conditions derived from spectral theory: Discrete systems with periodic feedback. SIAM Journal on Control, 10(1):1– 13, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Davis:1973:EOP

[Dav73] M. H. A. Davis. On the existence of optimal policies in stochastic control. SIAM Journal on Control, 11 (4):587–594, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Desoer:1972:CFS

[DC72] C. A. Desoer and F. M. Callier. Convolution feedback systems. SIAM Journal on Control, 10(4):737–746, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

deKorvin:1968:ATS

[de 68] Andre de Korvin. Approximation theorems on some classes of automata. SIAM Journal on Control, 6(3):422–436, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

DeJulio:1970:NSD

[De 70] S. De Julio. Numerical solution of dynamical optimization problems. SIAM Journal on Control, 8(2):135– 147, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

DeSantis:1974:CSC

[De 74] Romano M. De Santis. Causality, strict causality and invertibility for systems in Hilbert resolution spaces. SIAM Journal on Control, 12(3):536–553, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Demyanov:1968:SSO

[Dem68] V. F. Dem’yanov. The solution of some optimal control problems. SIAM Journal on Control, 6(1):59– 72, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

DeSantis:1973:SRC

[DeS73] Romano M. DeSantis. On state realization and causality decomposition for nonlinear systems. SIAM Journal on Control, 11(4):551–562, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

DAlessandro:1973:NAT

[DIR73] P. D’Alessandro, A. Isidori, and A. Ruberti. A new approach to the theory of canonical decomposition of linear dynamical systems. SIAM Journal on Control, 11(1):148–158, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

DAlessandro:1974:RST

[DIR74] Paolo D’Alessandro, Alberto Isidori, and Antonio Ruberti. Realization and structure theory of bilinear dynamical systems. SIAM Journal on Control, 12(3):517–535, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Davison:1970:SSC

[DK70] E. J. Davison and E. G. Kunze. Some sufficient conditions for the global and local controllability of nonlinear time-varying systems. SIAM Journal on Control, 8(4):489–497, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Delfour:1969:RPS

[DM69] M. C. Delfour and S. K. Mitter. Reachability of perturbed systems and min sup problems. SIAM Journal on Control, 7(4):521–533, ???? 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Delfour:1972:COI

[DM72a] M. C. Delfour and S. K. Mitter. Controllability and observability for infinite-dimensional systems. SIAM Journal on Control, 10(2):329–333, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Delfour:1972:COO

[DM72b] M. C. Delfour and S. K. Mitter. Controllability, observability and optimal feedback control of affine hereditary differential systems. SIAM Journal on Control, 10(2): 298–328, ???? 1972. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Delfour:1975:SIT

[DMM75] M. C. Delfour, C. McCalla, and S. K. Mitter. Stability and the infinite-time quadratic cost problem for linear hereditary differential systems. SIAM Journal on Control, 13 (1):48–88, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Dolezal:1973:EDM

[Dol73] Václav Doležal. Equations describing multidimensional causal systems. SIAM Journal on Control, 11(2): 306–322, ???? 1973. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Dolezal:1974:HN

[Dol74] Vaclav Dolezal. Hilbert networks.

I. SIAM Journal on Control, 12(4): 755–778, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Dooge:1964:ALS

[Doo64] James C. I. Dooge. Analysis of linear systems by means of Laguerre functions. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):396–408, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Dutkevich:1972:GLL

[DP72] Yu. G. Dutkevich and L. A. Petrosyan. Games with a “life-line”. The case of l-capture. SIAM Journal on Control, 10(1):40–47, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Demyanov:1967:MSC

[DR67] V. F. Dem’yanov and A. M. Rubinov. The minimization of a smooth convex functional on a convex set. SIAM Journal on Control, 5(2):280– 294, ???? 1967. CODEN SJCOA9.

ISSN 0036-1402 (print), 2469-4231 (electronic).

Demyanov:1968:MFN

[DR68] V. F. Dem’yanov and A. M. Rubinov. Minimization of functionals in normed spaces. SIAM Journal on Control, 6(1):73–88, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Dreyfus:1964:STO

[Dre64] Stuart E. Dreyfus. Some types of optimal control of stochastic systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1):120–134, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Das:1971:OCM

[DS71] P. C. Das and R. R. Sharma. On optimal controls for measure delaydifferential equations. SIAM Journal on Control, 9(1):43–61, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Duncan:1971:SSC

[DV71] Tyrone Duncan and Pravin Varaiya. On the solutions of a stochastic control system. SIAM Journal on Control, 9(3):354–371, ???? 1971. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Davis:1973:DPC

[DV73] M. H. A. Davis and P. Varaiya. Dynamic programming conditions for partially observable stochastic systems. SIAM Journal on Control, 11 (2):226–261, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Duncan:1975:SSC

[DV75] Tyrone Duncan and Pravin Varaiya.

On the solutions of a stochastic control system. II. SIAM Journal on Control, 13(5):1077–1092, August 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Dolezal:1975:HNI

[DZ75] Vaclav Dolezal and Armen Zemanian. Hilbert networks. II: Some qualitative properties. SIAM Journal on Control, 13(1):153–161, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Elliott:1975:NGP

[EF75] Robert J. Elliott and Avner Friedman. A note on generalized pursuitevasion games. SIAM Journal on Control, 13(1):105–109, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Evans:1974:ETP

[EG74] J. P. Evans and F. J. Gould. An existence theorem for penalty function theory. SIAM Journal on Control, 12(3):509–516, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Egorov:1966:OPD

[Ego66] A. I. Egorov. Optimal processes in distributed parameter systems and certain problems in invariance theory. SIAM Journal on Control, 4 (4):601–661, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Egorov:1967:NOC

[Ego67] A. I. Egorov. Necessary optimality conditions for distributed-parameter systems. SIAM Journal on Control, 5(3):352–408, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ewing:1964:OPA

[EH64] G. M. Ewing and W. R. Haseltine. Optimal programs for an ascending missile. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1):66–88, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Elliott:1973:SPL

[EKM73] R. J. Elliott, N. J. Kalton, and L. Markus. Saddle points for linear differential games. SIAM Journal on Control, 11(1):100–112, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Eaves:1971:GCP

[EZ71] B. Curtis Eaves and W. I. Zangwill. Generalized cutting plane algorithms. SIAM Journal on Control, 9(4):529–542, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Falk:1969:LMN

[Fal69] James E. Falk. Lagrange multipliers and nonconvex programs. SIAM Journal on Control, 7(4):534–545, ???? 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Fattorini:1964:TOC

[Fat64] H. O. Fattorini. Time-optimal control of solutions of operational differential equations. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1): 54–59, ???? 1964. CODEN JSM-CAL. ISSN 0887-4603.

Fattorini:1966:SRC

[Fat66] H. O. Fattorini. Some remarks on complete controllability. SIAM Journal on Control, 4(4):686–694, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Fattorini:1968:BCS

[Fat68a] H. O. Fattorini. Boundary control systems. SIAM Journal on Control, 6(3):349–385, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Fattorini:1968:RBB

[Fat68b] H. O. Fattorini. A remark on the “bang–bang” principle for linear control systems in infinite-dimensional space. SIAM Journal on Control, 6(1):109–113, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Fattorini:1975:BCT

[Fat75] H. O. Fattorini. Boundary control of temperature distributions in a parallelepipedon. SIAM Journal on Control, 13(1):1–13, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Faure:1972:ECC

[Fau72] Robert Faure. Sur l’existence et le comportement de cycles de certaines Équations Différentielles non Linéaires (solutions périodiques de systèmes autonomes). SIAM Journal on Control, 10(1):169–193, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Falb:1969:GTT

[FF69] P. L. Falb and M. I. Freedman. A generalized transform theory for causal operators. SIAM Journal on Control, 7(3):452–471, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Freedman:1969:NCA

[FFA69] M. I. Freedman, P. L. Falb, and J. Anton. A note on causality and analyticity. SIAM Journal on Control, 7(3):472–478, August 1969. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Freedman:1969:HSS

[FFZ69] M. I. Freedman, P. L. Falb, and G. Zames. A Hilbert space stability theory over locally compact Abelian groups. SIAM Journal on Control, 7 (3):479–495, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Funk:1970:SSC

[FG70] J. E. Funk and E. G. Gilbert. Some sufficient conditions for optimality in control problems with state space constraints. SIAM Journal on Control, 8(4):498–504, ???? 1970. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Freedman:1971:TFS

[FG71] M. I. Freedman and R. Glassey.

Tracking via feedback for systems with irrational transfer function. SIAM Journal on Control, 9(3):317– 338, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Filippov:1962:CQT

[Fil62] A. F. Filippov. On certain questions in the theory of optimal control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(1):76–84, ???? 1962. CODEN JSMCAL. ISSN 0887-4603.

Filippov:1967:CSD

[Fil67] A. F. Filippov. Classical solutions of differential equations with multivalued right-hand side. SIAM Journal on Control, 5(4):609–621, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). Translation from the Russian original in Vestnik Moskov. Univ., Ser. I 22, No. 3, 16–26 (1967).

Ford:1968:ISC

[FJ68] D. A. Ford and C. D. Johnson. Invariant subspaces and the controllability and observability of linear dynamical systems. SIAM Journal on Control, 6(4):553–558, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Fleming:1968:OCP

[Fle68] Wendell H. Fleming. Optimal control of partially observable diffusions. SIAM Journal on Control, 6(2):194– 214, ???? 1968. CODEN SJCOA9.

ISSN 0036-1402 (print), 2469-4231 (electronic).

Fleming:1971:SCS

[Fle71] Wendell H. Fleming. Stochastic control for small noise intensities. SIAM Journal on Control, 9(3):473–517, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Flynn:1973:LMB

[Fly73] James O. Flynn. Lion and man: The boundary constraint. SIAM Journal on Control, 11(3):397–411, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Flynn:1974:LMG

[Fly74a] James Flynn. Lion and man: The general case. SIAM Journal on Control, 12(4):581–597, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Flynn:1974:SRM

[Fly74b] James Flynn. Some results on maxmin pursuit. SIAM Journal on Control, 12(1):53–69, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Formalskii:1967:CRC

[For67] A. M. Formal′skiĭ. Construction of the region of controllability for systems with bounded-impulse controls. SIAM Journal on Control, 5(2):211– 221, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Forney:1975:MBR

[For75] G. David Forney, Jr. Minimal bases of rational vector spaces, with applications to multivariable linear systems. SIAM Journal on Control, 13(3):493–520, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Freedman:1968:STV

[Fre68] M. I. Freedman. L_{2}-stability of time-varying systems—construction of multipliers with prescribed phase characteristics. SIAM Journal on Control, 6(4):559–578, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Freedman:1972:PFN

[Fre72] Marvin I. Freedman. Phase function norm estimates for stability of systems with monotone nonlinearities. SIAM Journal on Control, 10(1):99– 111, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Fabian:1974:GSD

[FW74] E. Fabian and W. M. Wonham.

Generic solvability of the decoupling problem. SIAM Journal on Control, 12(4):688–694, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Fujisawa:1967:IPS

[FY67] Toshio Fujisawa and Yutaka Yasuda.

An iterative procedure for solving the time-optimal regulator problem. SIAM Journal on Control, 5(4):501– 512, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Freedman:1968:LVC

[FZ68] M. Freedman and G. Zames. Logarithmic variation criteria for the stability of systems with time-varying gains. SIAM Journal on Control, 6 (3):487–507, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

GalChuk:1969:OCS

[Gal69] L. I. Gal’Chuk. Optimal control of systems described by parabolic equations. SIAM Journal on Control, 7(4):546–558, November 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gambill:1963:GCE

[Gam63] R. A. Gambill. Generalized curves and the existence of optimal controls. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3):246–260, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Gamkrelidze:1965:SEP

[Gam65] R. V. Gamkrelidze. On some extremal problems in the theory of differential equations with applications to the theory of optimal control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):106–128, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Garfinkel:1963:SGP

[Gar63] Boris Garfinkel. A solution of the Goddard problem. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3):349–368, ???? 1963. CODEN JSMCAL. ISSN 0887-4603. See errata [Gar64].

Garfinkel:1964:ESG

[Gar64] Boris Garfinkel. Errata: A solution of the Goddard problem. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1):135, ???? 1964. CO-DEN JSMCAL. ISSN 0887-4603. See [Gar63].

Gehner:1974:NSO

[Geh74] Kenneth R. Gehner. Necessary and sufficient optimality conditions for the Fritz John problem with linear equality constraints. SIAM Journal on Control, 12(1):140–149, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gilbert:1963:COM

[Gil63] Elmer G. Gilbert. Controllability and observability in multivariable control systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(2): 128–151, ???? 1963. CODEN JSM-CAL. ISSN 0887-4603.

Gilbert:1966:IPC

[Gil66] Elmer G. Gilbert. An iterative procedure for computing the minimum of a quadratic form on a convex set. SIAM Journal on Control, 4(1):61– 80, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gilbert:1967:NFP

[Gil67] Elmer G. Gilbert. A note on the fixed-point method of J. E. Rubio.

SIAM Journal on Control, 5(4):513– 514, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gilbert:1969:DMS

[Gil69] Elmer G. Gilbert. The decoupling of multivariable systems by state feedback. SIAM Journal on Control, 7 (1):50–63, February 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gindes:1967:OCL

[Gin67a] V. B. Gindes. Optimal control in a linear system under conflict. SIAM Journal on Control, 5(1):163–169, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gindes:1967:POJ

[Gin67b] V. B. Gindes. A problem of optimal joint control. SIAM Journal on Control, 5(2):222–227, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Girsanov:1969:CRB

[Gir69] I. V. Girsanov. Certain relations between the Bellman and Krotov functions for dynamic programming problems. SIAM Journal on Control, 7(1):64–67, February 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gabasov:1972:HON

[GK72] R. Gabasov and F. M. Kirillova. High order necessary conditions for optimality. SIAM Journal on Control, 10(1):127–168, ???? 1972. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gamkrelidze:1974:DGE

[GK74] R. V. Gamkrelidze and G. L.

Kharatishvili. A differential game of evasion with nonlinear control. SIAM Journal on Control, 12(2): 332–349, May 1974. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Glover:1969:IPF

[Glo69] Fred Glover. Integer programming over a finite additive group. SIAM Journal on Control, 7(2):213–231, ???? 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Goh:1966:NCS

[Goh66a] B. S. Goh. Necessary conditions for singular extremals involving multiple control variables. SIAM Journal on Control, 4(4):716–731, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Goh:1966:SVS

[Goh66b] B. S. Goh. The second variation for the singular Bolza problem. SIAM Journal on Control, 4(2):309–325, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Goldstein:1965:CPO

[Gol65a] A. A. Goldstein. Convex programming and optimal control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):142–146, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Goldstein:1965:SD

[Gol65b] A. A. Goldstein. On steepest descent. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):147–151, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Goldstein:1966:MFN

[Gol66] A. A. Goldstein. Minimizing functionals on normed-linear spaces. SIAM Journal on Control, 4(1):81– 89, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gordon:1972:CST

[Gor72a] S. P. Gordon. On converses to the stability theorems for difference equations. SIAM Journal on Control, 10(1):76–81, ???? 1972. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gordon:1972:STP

[Gor72b] S. P. Gordon. A stability theory for perturbed difference equations. SIAM Journal on Control, 10(4): 671–678, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Grigoriadis:1969:PMS

[GR69] M. D. Grigoriadis and K. Ritter. A parametric method for semidefinite quadratic programs. SIAM Journal on Control, 7(4):559–577, November 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Graham:1975:BVC

[GR75] Keith D. Graham and David L. Russell. Boundary value control of the wave equation in a spherical region. SIAM Journal on Control, 13(1): 174–196, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Grimmell:1967:EPC

[Gri67] W. C. Grimmell. The existence of piecewise continuous fuel optimal controls. SIAM Journal on Control, 5(4):515–519, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ginsburg:1966:FTP

[GS66] Seymour Ginsburg and Edwin H.

Spanier. Finite-turn pushdown automata. SIAM Journal on Control, 4(3):429–453, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Goldwyn:1967:TOC

[GSG67] R. M. Goldwyn, K. P. Sriram, and M. Graham. Time optimal control of a linear diffusion process. SIAM Journal on Control, 5(2):295–308, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Goldstein:1973:KES

[GT73] A. A. Goldstein and R. A. Tapia.

Kantorovich estimates for some convex programs. SIAM Journal on Control, 11(2):375–377, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Guinn:1965:PBS

[Gui65a] T. Guinn. The problem of bounded space coordinates as a problem of Hestenes. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2):181–190, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Guinn:1965:WHV

[Gui65b] T. Guinn. Weakened hypotheses for the variational problem considered by Hestenes. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3): 418–423, ???? 1965. CODEN JSM-CAL. ISSN 0887-4603.

Guignard:1969:GKT

[Gui69] Monique Guignard. Generalized Kuhn–Tucker conditions for mathematical programming problems in a Banach space. SIAM Journal on Control, 7(2):232–241, ???? 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Gurley:1964:OTT

[Gur64] Joseph G. Gurley. Optimal-thrust trajectories in an arbitrary gravitational field. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):423–432, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Haas:1975:NCP

[Haa75] Violet B. Haas. On normality and conjugate point criteria for singular extremals. SIAM Journal on Control, 13(6):1172–1182, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Hack:1973:ECR

[Hac73] Thomas G. Hack. Extensions of the concept of regular synthesis as a sufficient condition for optimality. SIAM Journal on Control, 11 (2):358–374, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Haddad:1973:CBI

[Had73] Emile K. Haddad. A criterion for the bounded-input, bounded-output stability of time-varying nonlinear systems. SIAM Journal on Control, 11 (2):286–305, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Hajek:1971:GTT

[Háj71] Otomar Hájek. Geometric theory of time-optimal control. SIAM Journal on Control, 9(3):339–350, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Halkin:1964:GLB

[Hal64] Hubert Halkin. A generalization of LaSalle’s “bang–bang” principle. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):199–202, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Halkin:1966:MPP

[Hal66] Hubert Halkin. A maximum principle of the Pontryagin type for systems described by nonlinear difference equations. SIAM Journal on Control, 4(1):90–111, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Halanay:1968:DGD

[Hal68a] A. Halanay. Differential games with delay. SIAM Journal on Control, 6 (4):579–593, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Halanay:1968:OCS

[Hal68b] A. Halanay. Optimal controls for systems with time lag. SIAM Journal on Control, 6(2):215–234, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Hal71].

Halanay:1971:EOC

[Hal71] A. Halanay. Erratum: “Optimal controls for systems with time lag”. SIAM Journal on Control, 9(1):142, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Hal68b].

Halkin:1974:IFO

[Hal74] Hubert Halkin. Implicit functions and optimization problems without continuous differentiability of the data. SIAM Journal on Control, 12(2):229–236, May 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Harvey:1964:MFR

[Har64] C. A. Harvey. Modes of finite response time control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2 (1):60–65, ???? 1964. CODEN JSM-CAL. ISSN 0887-4603.

Haseltine:1968:EST

[Has68] W. R. Haseltine. Existence and stability theorems for exterior ballistics. SIAM Journal on Control, 6(3):386– 400, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Haussmann:1971:OSC

[Hau71] U. G. Haussmann. Optimal stationary control with state control dependent noise. SIAM Journal on Control, 9(2):184–198, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Haussmann:1973:SLS

[Hau73a] U. G. Haussmann. Stability of linear systems with control dependent noise. SIAM Journal on Control, 11 (2):382–394, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Hautus:1973:NCM

[Hau73b] M. L. J. Hautus. Necessary conditions for multiple constraint optimization problems. SIAM Journal on Control, 11(4):653–669, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Haussmann:1974:EMS

[Hau74] U. G. Haussmann. On the existence of moments of stationary linear systems with multiplicative noise. SIAM Journal on Control, 12(1): 99–105, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Haynes:1966:OTS

[Hay66] George W. Haynes. On the optimality of a totally singular vector control: an extension of the Green’s theorem approach to higher dimensions. SIAM Journal on Control, 4(4):662– 677, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ho:1963:COC

[HB63] Yu-Chi Ho and Piero B. Brentani. On computing optimal control with inequality constraints. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3):319–348, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Hermes:1964:CSP

[Her64] H. Hermes. Controllability and the singular problem. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):241–260, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Hermes:1967:CCA

[Her67] H. Hermes. On the closure and convexity of attainable sets in finite and infinite dimensions. SIAM Journal on Control, 5(3):409–417, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See errata [Her68].

Hermes:1968:ECC

[Her68] H. Hermes. Errata: On the closure and convexity of attainable sets in finite and infinite dimensions. SIAM Journal on Control, 6(4):594–595, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Her67].

Hermes:1972:GTO

[Her72] Henry Hermes. The geometry of time-optimal control. SIAM Journal on Control, 10(2):221–229, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Hermes:1974:LGC

[Her74] H. Hermes. On local and global controllability. SIAM Journal on Control, 12(2):252–261, May 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Hestenes:1965:VTO

[Hes65] Magnus R. Hestenes. On variational theory and optimal control theory. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):23–48, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Hewer:1975:PDM

[Hew75] G. A. Hewer. Periodicity, detectability and the matrix Riccati equation. SIAM Journal on Control, 13(6): 1235–1251, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Heymann:1972:PSL

[Hey72] Michael Heymann. The prime structure of linear dynamical systems. SIAM Journal on Control, 10(3): 460–469, ???? 1972. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Hermes:1963:NCP

[HH63] H. Hermes and G. Haynes. On the nonlinear control problem with control appearing linearly. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(2):85–108, ???? 1963. CO-DEN JSMCAL. ISSN 0887-4603.

Holtzman:1966:DCM

[HH66] J. M. Holtzman and H. Halkin. Discretional convexity and the maximum principle for discrete systems. SIAM Journal on Control, 4(2):263– 275, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Haynes:1970:NCL

[HH70] G. W. Haynes and H. Hermes. Nonlinear controllability via Lie theory. SIAM Journal on Control, 8(4):450– 460, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ho:1966:CIP

[HK66] Yu-Chi Ho and R. L. Kashyap. A class of iterative procedures for linear inequalities. SIAM Journal on Control, 4(1):112–115, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ho:1966:OTM

[Ho66] Yu-Chi Ho. Optimal terminal maneuver and evasion strategy. SIAM Journal on Control, 4(3):421–428, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Hautus:1973:UTL

[HO73] M. L. J. Hautus and G. J. Olsder. A uniqueness theorem for linear control systems with coinciding reachable sets. SIAM Journal on Control, 11(3):412–416, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Hofri:1975:SSD

[Hof75] Micha Hofri. Stochastic scheduling design in multiserver systems. SIAM Journal on Control, 13(3):736–748, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Holtzman:1968:ASL

[Hol68] J. M. Holtzman. Analysis of statistical linearization of nonlinear control systems. SIAM Journal on Control, 6(2):235–243, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Holland:1974:SNO

[Hol74] Charles J. Holland. Small noise open loop control. SIAM Journal on Control, 12(3):380–388, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Holland:1975:GOL

[Hol75] Charles J. Holland. Gaussian open loop control problems. SIAM Journal on Control, 13(3):545–551, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Howe:1973:NCE

[How73] Stephen Howe. New conditions for exactness of a simple penalty function. SIAM Journal on Control, 11 (2):378–381, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Heymann:1970:TEL

[HT70] Michael Heymann and John A.

Thorpe. Transfer equivalence of linear dynamical systems. SIAM Journal on Control, 8(1):19–40, February 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Huang:1972:OCP

[Hua72] Sheng-Chao Huang. Optimal control problems with a system of integral equations and restricted phase coordinates. SIAM Journal on Control, 10(1):14–36, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ikeda:1972:SLS

[IMK72] Masao Ikeda, Hajime Maeda, and Shinzo Kodama. Stabilization of linear systems. SIAM Journal on Control, 10(4):716–729, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ikeda:1975:EFL

[IMK75] Masao Ikeda, Hajime Maeda, and Shinzo Kodama. Estimation and feedback in linear time-varying systems: a deterministic theory. SIAM Journal on Control, 13(2):304–326, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ikeda:1975:CSP

[IS75] Mineo Ikeda and Kentaro Sakamoto. On the concept of symmetry in Pontryagin’s maximum principle. SIAM Journal on Control, 13(2):389–399, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Inaba:1975:GRP

[IT75] Hiroshi Inaba and Byron D. Tapley. Generalized random processes: a theory and the white Gaussian process. SIAM Journal on Control, 13 (3):719–735, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jacobs:1967:RSR

[Jac67a] Marc Q. Jacobs. Remarks on some recent extensions of Filippov’s implicit functions lemma. SIAM Journal on Control, 5(4):622–627, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jacobs:1967:SET

[Jac67b] Marc Q. Jacobs. Some existence theorems for linear optimal control problems. SIAM Journal on Control, 5(3):418–437, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jacobs:1969:AIM

[Jac69a] Marc Q. Jacobs. On the approximation of integrals of multivalued functions. SIAM Journal on Control, 7 (1):158–177, ???? 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jacobson:1969:NNC

[Jac69b] D. H. Jacobson. A new necessary condition of optimality for singular control problems. SIAM Journal on Control, 7(4):578–595, November 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jacobson:1970:SCN

[Jac70] D. H. Jacobson. Sufficient conditions for nonnegativity of the second variation in singular and nonsingular control problems. SIAM Journal on Control, 8(3):403–423, August 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jameson:1973:IPL

[JK73] Antony Jameson and Eliezer Kreindler. Inverse problem of linear optimal control. SIAM Journal on Control, 11(1):1–19 (or 1–9??), ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jones:1970:GMB

[JM70] Arnold P. Jones and Garth P. Mc-Cormick. A generalization of the method of Balakrishnan. inequality constraints and initial conditions. SIAM Journal on Control, 8(2):218– 225, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Johnson:1973:MET

[Joh73] Timothy L. Johnson. Minimumenergy terminal state control of first order linear hyperbolic systems in one spatial variable using the method of characteristics. SIAM Journal on Control, 11(1):119–129, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jordan:1964:OCA

[JP64] B. W. Jordan and E. Polak. Optimal control of aperiodic discretetime systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3): 332–346, ???? 1964. CODEN JSM-CAL. ISSN 0887-4603.

Jurdjevic:1970:ACS

[Jur70] Velimir Jurdjevic. Abstract control systems. controllability and observability. SIAM Journal on Control, 8 (3):424–439, August 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Jurdjevic:1972:CCP

[Jur72] Velimir Jurdjevic. Certain controllability properties of analytic control systems. SIAM Journal on Control, 10(2):354–360, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kailath:1969:ARI

[Kai69] Thomas Kailath. Application of a resolvent identity to a linear smoothing problem. SIAM Journal on Control, 7(1):68–74, ???? 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kailath:1972:NLS

[Kai72] Thomas Kailath. A note on least squares estimation by the innovations method. SIAM Journal on Control, 10(3):477–486, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kakumanu:1972:NCT

[Kak72] Prasadarao Kakumanu. Nondiscounted continuous time Markovian decision process with countable state space. SIAM Journal on Control, 10 (1):210–220, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kalman:1963:MDL

[Kal63] R. E. Kalman. Mathematical description of linear dynamical systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(2):152–192, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Kushner:1970:CLF

[KB70] Harold J. Kushner and Daniel I. Barnea. On the control of a linear functional-differential equation with quadratic cost. SIAM Journal on Control, 8(2):257–272, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Karush:1966:OPS

[KD66] W. Karush and R. E. Dear. An optimal procedure for an N-stage learning process. SIAM Journal on Control, 4(1):116–129, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kelley:1964:TAS

[Kel64] Henry J. Kelley. A transformation approach to singular subarcs in optimal trajectory and control problems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):234–240, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Krasnakevich:1969:GPC

[KH69] J. R. Krasnakevich and R. A. Haddad. Generalized predictioncorrection estimation. SIAM Journal on Control, 7(3):496–511, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See comments [Wei77].

Kirillova:1963:CFO

[Kir63] F. M. Kirillova. On the correctness of the formulation of an optimal control problem. Journal of the Society for Industrial and Applied Mathematics.

Series A, On control, 1(2):224–239, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Kirillova:1967:AFA

[Kir67] F. M. Kirillova. Applications of functional analysis to the theory of optimal processes. SIAM Journal on Control, 5(1):25–50, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Klessig:1974:GTC

[Kle74] R. Klessig. General theory of convergence for constrained optimization algorithms that use antizigzagging provisions. SIAM Journal on Control, 12(4):598–608, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Knowles:1975:LVM

[Kno75] Gregory Knowles. Lyapunov vector measures. SIAM Journal on Control, 13(2):294–303, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kall:1975:MTS

[KO75] P. Kall and W. Oettli. Measurability theorems for stochastic extremals. SIAM Journal on Control, 13(5):994–998, August 1975. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Komkov:1968:OCT

[Kom68] V. Komkov. The optimal control of a transverse vibration of a beam. SIAM Journal on Control, 6(3):401– 421, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Komkov:1970:OCV

[Kom70] Vadim Komkov. Optimal control of vibrating thin plates. SIAM Journal on Control, 8(2):273–304, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Klessig:1972:EIP

[KP72] R. Klessig and E. Polak. Efficient implementations of the Polak–Ribière conjugate gradient algorithm. SIAM Journal on Control, 10(3):524–549, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Klessig:1973:APG

[KP73] R. Klessig and E. Polak. An adaptive precision gradient method for optimal control. SIAM Journal on Control, 11(1):80–93, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kirchgassner:1966:SPN

[KR66] K. Kirchgässner and K. Ritter. On stationary points of nonlinear maximum-problems in Banach spaces. SIAM Journal on Control, 4(4):732–739, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Krall:1964:ATL

[Kra64a] Allan M. Krall. The asymptotes of the time lag root-locus. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):370–372, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Krall:1964:SCF

[Kra64b] Allan M. Krall. Stability criteria for feedback systems with a time lag. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):160–170, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Kreindler:1969:SCL

[Kre69] Eliezer Kreindler. On sensitivity of closed loop nonlinear optimal control systems. SIAM Journal on Control, 7(3):512–520, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Krener:1973:ECS

[Kre73] Arthur J. Krener. On the equivalence of control systems and the linearization of nonlinear systems. SIAM Journal on Control, 11(4): 670–676, ???? 1973. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Krener:1974:GCT

[Kre74] Arthur J. Krener. A generalization of Chow’s theorem and the bang–bang theorem to nonlinear control problems. SIAM Journal on Control, 12 (1):43–52, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Krener:1975:BNR

[Kre75] Arthur J. Krener. Bilinear and nonlinear realizations of input-output maps. SIAM Journal on Control, 13(4):827–834, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Krolak:1968:FEF

[Kro68] Patrick D. Krolak. Further extensions of Fibonaccian search to nonlinear programming problems. SIAM Journal on Control, 6(2):258–265, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kullback:1966:ITD

[Kul66] S. Kullback. An informationtheoretic derivation of certain limit relations for a stationary Markov chain. SIAM Journal on Control, 4 (3):454–459, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kullback:1967:EIT

[Kul67] S. Kullback. An extension of an information-theoretic derivation of certain limit relations for a Markov chain. SIAM Journal on Control, 5(1):51–53, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kunzi:1966:DMN

[Kün66] H. P. Künzi. The duoplex method in nonlinear programming. SIAM Journal on Control, 4(1):130–138, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kushner:1964:DES

[Kus64] Harold J. Kushner. On the differential equations satisfied by conditional probability densities of Markov processes, with applications. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1):106–119, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Kushner:1965:EOS

[Kus65a] Harold J. Kushner. On the existence of optimal stochastic controls. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3):463–474, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Kushner:1965:SCO

[Kus65b] Harold J. Kushner. Sufficient conditions for the optimality of a stochastic control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3):499–508, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Kushner:1967:CTS

[Kus67a] H. Kushner. Converse theorems for stochastic Liapunov functions. SIAM Journal on Control, 5(2):228– 233, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kushner:1967:ODS

[Kus67b] H. J. Kushner. Optimal discounted stochastic control for diffusion processes. SIAM Journal on Control, 5(4):520–531, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kushner:1968:OCS

[Kus68] H. J. Kushner. On the optimal control of a system governed by a linear parabolic equation with white noise inputs. SIAM Journal on Control, 6(4):596–614, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kushner:1970:FLD

[Kus70] Harold J. Kushner. Filtering for linear distributed parameter systems. SIAM Journal on Control, 8(3):346– 359, August 1970. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kushner:1972:NCC

[Kus72] H. J. Kushner. Necessary conditions for continuous parameter stochastic optimization problems. SIAM Journal on Control, 10(3):550–565, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Kwakernaak:1975:PLD

[Kwa75] Huibert Kwakernaak. Periodic linear differential stochastic processes. SIAM Journal on Control, 13(2): 400–413, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

LaSalle:1962:SC

[LaS62] J. P. LaSalle. Stability and control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(1):3–15, ???? 1962. CODEN JSMCAL. ISSN 0887-4603.

Lee:1972:CDT

[LCB72] Kwang Yun Lee, Shui-Nee Chow, and Robert O. Barr. On the control of discrete-time distributed parameter systems. SIAM Journal on Control, 10(2):361–376, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lee:1963:SCT

[Lee63] E. B. Lee. A sufficient condition in the theory of optimal control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3):241–245, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Lefschetz:1965:LSC

[Lef65] Solomon Lefschetz. Liapunov stability and controls. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1): 1–6, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Lehnigk:1967:LDM

[Leh67] Siegfried H. Lehnigk. Liapunov’s direct method and the number of zeros with positive real parts of a polynomial with constant complex coefficients. SIAM Journal on Control, 5(2):234–244, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Leitmann:1965:SGA

[Lei65] G. Leitmann. Some geometrical aspects of optimal processes. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):53–65, ???? 1965. CO-DEN JSMCAL. ISSN 0887-4603.

Lellouche:1970:FCO

[Lel70] Gerald S. Lellouche. A frequency criterion for oscillatory solutions. SIAM Journal on Control, 8(2):202– 206, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lee:1969:DMS

[LH69] Muo S. Lee and C. S. Hsu. On the τ-decomposition method of stability analysis for retarded dynamical systems. SIAM Journal on Control, 7(2):242–259, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lindquist:1973:FCL

[Lin73] Anders Lindquist. On feedback control of linear stochastic systems. SIAM Journal on Control, 11(2): 323–343, ???? 1973. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lindquist:1974:NAO

[Lin74a] Anders Lindquist. New algorithm for optimal filtering of discrete-time stationary processes. SIAM Journal on Control, 12(4):736–746, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lindquist:1974:OFC

[Lin74b] Anders Lindquist. Optimal filtering of continuous-time stationary processes by means of the backward innovation process. SIAM Journal on Control, 12(4):747–754, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Liu:1966:EEC

[LL66] Ruey-Wen Liu and R. Jeffrey Leake. Exhaustive equivalence classes of optimal systems with separable controls. SIAM Journal on Control, 4 (4):678–685, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lo:1975:GBS

[Lo75] James Ting-Ho Lo. Global bilinearization of systems with control appearing linearly. SIAM Journal on Control, 13(4):879–885, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lobry:1970:CSN

[Lob70] Claude Lobry. Contrôlabilité des systèmes non linéaires. (French) [Controllability of nonlinear systems]. SIAM Journal on Control, 8(4):573–605, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Lob76].

Lobry:1974:CNS

[Lob74] C. Lobry. Controllability of nonlinear systems on compact manifolds. SIAM Journal on Control, 12(1):1– 4, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lobry:1976:ECS

[Lob76] C. Lobry. Erratum: Contrôlabilité des systèmes non linéaires. (French) [Erratum: Controllability of nonlinear systems]. SIAM Journal on Control and Optimization, 14(2):387, February 1976. CODEN SJCODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [Lob70].

Loridan:1973:POUa

[Lor73a] Pierre Loridan. Sur un procédé d’optimisation utilisant simultanément les méthodes de pénalisation et des variations locales. I. SIAM Journal on Control, 11(1):159–172, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Loridan:1973:POUb

[Lor73b] Pierre Loridan. Sur un procédé d’optimization utilisant simultanement les méthodes de pénalisation et des variations locales. II. (French) [On an optimization procedure using simultaneously penalty and local variation methods]. SIAM Journal on Control, 11(1):173–184, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lukes:1969:QCD

[LR69] D. L. Lukes and D. L. Russell. The quadratic criterion for distributed systems. SIAM Journal on Control, 7(1):101–121, February 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Levitt:1975:CMT

[LS75] Norman Levitt and Héctor J. Sussmann. On controllability by means of two vector fields. SIAM Journal on Control, 13(6):1271–1281, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lukes:1969:ORN

[Luk69] D. L. Lukes. Optimal regulation of nonlinear dynamical systems. SIAM Journal on Control, 7(1):75–100, February 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lukes:1971:EFC

[Luk71] D. L. Lukes. Equilibrium feedback control in linear games with quadratic costs. SIAM Journal on Control, 9(2):234–252, ???? 1971.

CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Luk72a].

Lukes:1972:EEF

[Luk72a] D. L. Lukes. Erratum: Equilibrium feedback control in linear games with quadratic costs. SIAM Journal on Control, 10(3):566, ???? 1972. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Luk71].

Lukes:1972:GCN

[Luk72b] D. L. Lukes. Global controllability of nonlinear systems. SIAM Journal on Control, 10(1):112–126, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Luk73].

Lukes:1973:EGC

[Luk73] D. L. Lukes. Erratum: “Global controllability of nonlinear systems” (SIAM J. Control 10 (1972), 112– 126). SIAM Journal on Control, 11(1):186, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Luk72b].

Lukes:1974:GCD

[Luk74] D. L. Lukes. Global controllability of disturbed nonlinear equations. SIAM Journal on Control, 12(4): 695–704, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Lo:1975:SCR

[LW75] James Ting Ho Lo and Alan S. Willsky. Stochastic control of rotational processes with one degree of freedom. SIAM Journal on Control, 13(4):886–898, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Leake:1967:CSC

[LwL67] R. J. Leake and Ruey wen Liu. Construction of suboptimal control sequences. SIAM Journal on Control, 5(1):54–63, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Martinet:1974:MDP

[MA74] B. Martinet and A. Auslender. Méthodes de décomposition pour la minimisation d’une fonction sur un espace produit (French) [Decomposition methods for function minimization in a product space]. SIAM Journal on Control, 12(4):635–642, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Maeda:1973:SCV

[Mae73] Hajime Maeda. Stability considerations for a Volterra integral equation with discontinuous nonlinearity. SIAM Journal on Control, 11 (2):202–214, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Malanowski:1969:OCV

[Mal69] Kazimierz Malanowski. On optimal control of the vibrating string. SIAM Journal on Control, 7(2):260– 271, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Mangasarian:1963:SCN

[Man63] O. L. Mangasarian. Stability criteria for nonlinear ordinary differential equations. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3):311–318, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Mangasarian:1965:PCF

[Man65] O. L. Mangasarian. Pseudo-convex functions. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2):281–290, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Mangasarian:1966:SCO

[Man66] O. L. Mangasarian. Sufficient conditions for the optimal control of nonlinear systems. SIAM Journal on Control, 4(1):139–152, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Mangasarian:1975:ULN

[Man75] O. L. Mangasarian. Unconstrained Lagrangians in nonlinear programming. SIAM Journal on Control, 13(4):772–791, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Markus:1965:CNP

[Mar65] Lawrence Markus. Controllability of nonlinear processes. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3 (1):78–90, ???? 1965. CODEN JSM-CAL. ISSN 0887-4603.

Maurer:1975:ECJ

[Mau75] Helmut Maurer. Example of a continuous junction for a singular control problem of even order. SIAM Journal on Control, 13(4):899–903, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

McKnight:1973:RGP

[MB73] R. S. McKnight and W. E. Bosarge, Jr. The Ritz–Galerkin procedure for parabolic control problems. SIAM Journal on Control, 11(3):510–524, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

McCormick:1975:AMN

[McC75] Garth P. McCormick. An arc method for nonlinear programming. SIAM Journal on Control, 13(6): 1194–1216, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

McGill:1965:OCI

[McG65] Robert McGill. Optimum control, inequality state constraints, and the generalized Newton–Raphson algorithm. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2):291–298, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

McMillan:1962:I

[McM62] Brockway McMillan. Introduction. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(1):1–2, ???? 1962. CO-DEN JSMCAL. ISSN 0887-4603.

McShane:1967:RCV

[McS67] E. J. McShane. Relaxed controls and variational problems. SIAM Journal on Control, 5(3):438–485, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Mitchell:1974:FPP

[MDM74] B. F. Mitchell, V. F. Dem’yanov, and V. N. Malozemov. Finding the point of a polyhedron closest to the origin. SIAM Journal on Control, 12 (1):19–26, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Meditch:1967:OPD

[Med67] James S. Meditch. Orthogonal projection and discrete optimal linear smoothing. SIAM Journal on Control, 5(1):74–89, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Merriam:1964:AIS

[Mer64] C. W. Merriam III. An algorithm for the iterative solution of a class of two-point boundary value problems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1):1–10, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Meyer:1965:ELF

[Mey65] K. R. Meyer. On the existence of Lyapunov function for the problem of Lur’e. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3):373–383, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Meyer:1970:FTC

[Mey70a] Gunter H. Meyer. On fixed time control problems in a Banach space. SIAM Journal on Control, 8(3):383– 395, August 1970. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Meyer:1970:VFO

[Mey70b] Robert Meyer. The validity of a family of optimization methods. SIAM Journal on Control, 8(1):41– 54, February 1970. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Meyer:1973:DMF

[Mey73] Gerard G. L. Meyer. A drivable method of feasible directions. SIAM Journal on Control, 11(1):113–118, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Meyer:1974:AFW

[Mey74] Gerard G. L. Meyer. Accelerated Frank–Wolfe algorithms. SIAM Journal on Control, 12(4):655–663, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Mitter:1971:CPA

[MF71] S. K. Mitter and R. Foulkes. Controllability and pole assignment for discrete time linear systems defined over arbitrary fields. SIAM Journal on Control, 9(1):1–7, February 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Mond:1968:DCP

[MH68] Bertram Mond and Morgan Hanson. Duality for control problems. SIAM Journal on Control, 6(1):114– 120, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Michel:1974:SAI

[Mic74] Anthony N. Michel. Stability analysis of interconnected systems. SIAM Journal on Control, 12(3):554–579, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Miele:1965:EPA

[Mie65] Angelo Miele. Extremal problems in aerodynamics. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1): 129–141, ???? 1965. CODEN JSM-CAL. ISSN 0887-4603.

Miller:1968:FSC

[Mil68] Bruce L. Miller. Finite state continuous time Markov decision processes with a finite planning horizon. SIAM Journal on Control, 6(2):266– 280, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Miller:1972:SDE

[Mil72] John J. H. Miller. On the stability of differential equations. SIAM Journal on Control, 10(4):639–648, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Mirica:1969:ASO

[Mir69] Stefan Mirică. On the admissible synthesis in optimal control theory and differential games. SIAM Journal on Control, 7(2):292–316, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Mishchenko:1965:CPP

[Mis65] E. F. Mishchenko. On a certain problem for parabolic differential equations connected with optimal pursuit. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):66–70, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Mishchenko:1974:PEE

[Mis74] E. F. Mishchenko. On the problem of evading the encounter in differential games. SIAM Journal on Control, 12(2):300–310, May 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Minamide:1972:LBP

[MN72] N. Minamide and K. Nakamura. Linear bounded phase coordinate control problems under certain regularity and normality conditions. SIAM Journal on Control, 10(1):82–92, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Makowski:1974:OCP

[MN74] Karol Makowski and Lucien W. Neustadt. Optimal control problems with mixed control-phase variable equality and inequality constraints. SIAM Journal on Control, 12(2):184–228, May 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Molinari:1973:ERA

[Mol73a] B. P. Molinari. Equivalence relations for the algebraic Riccati equation. SIAM Journal on Control, 11 (2):272–285, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Molinari:1973:SSA

[Mol73b] B. P. Molinari. The stabilizing solution of the algebraic Riccati equation. SIAM Journal on Control, 11 (2):262–271, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Molinari:1975:NQF

[Mol75] B. P. Molinari. Nonnegativity of a quadratic functional. SIAM Journal on Control, 13(4):792–806, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Moreau:1966:QPM

[Mor66] J. J. Moreau. Quadratic programming in mechanics: Dynamics of one-sided constraints. SIAM Journal on Control, 4(1):153–158, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Morse:1971:OCS

[Mor71] A. S. Morse. Output controllability and system synthesis. SIAM Journal on Control, 9(2):143–148, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Morse:1973:SIL

[Mor73] A. S. Morse. Structural invariants of linear multivariable systems. SIAM Journal on Control, 11(3):446–465, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Moyer:1973:SCS

[Moy73] H. Gardner Moyer. Sufficient conditions for a strong minimum in singular control problems. SIAM Journal on Control, 11(4):620–636, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Messerli:1969:SON

[MP69] E. J. Messerli and E. Polak. On second order necessary conditions of optimality. SIAM Journal on Control, 7(2):272–291, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

McDanell:1971:NCJ

[MP71a] J. P. McDanell and W. F. Powers. Necessary conditions joining optimal singular and nonsingular subarcs. SIAM Journal on Control, 9 (2):161–173, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Meyer:1971:AMS

[MP71b] Gerard G. L. Meyer and E. Polak. Abstract models for the synthesis of optimization algorithms. SIAM Journal on Control, 9(4):547–560, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Mangasarian:1971:DSM

[MS71] O. L. Mangasarian and L. L. Schumaker. Discrete splines via mathematical programming. SIAM Journal on Control, 9(2):174–183, ????

1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Morse:1973:SII

[MS73] A. S. Morse and L. M. Silverman.

Structure of index-invariant systems. SIAM Journal on Control, 11(2): 215–225, ???? 1973. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Moore:1975:ECM

[MS75] John B. Moore and S. Sankar Sengupta. Existence and control of Markov chains in systems of deterministic motion. SIAM Journal on Control, 13(5):1103–1114, August 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

McCormick:1972:GPM

[MT72] G. P. McCormick and R. A. Tapia.

The gradient projection method under mild differentiability conditions. SIAM Journal on Control, 10(1):93– 98, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Murty:1969:TTS

[Mur69] Katta G. Murty. On the tours of a traveling salesman. SIAM Journal on Control, 7(1):122–131, February 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Murphy:1974:CEP

[Mur74] Frederic H. Murphy. A class of exponential penalty functions. SIAM Journal on Control, 12(4):679–687, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Morse:1970:DPA

[MW70] A. S. Morse and W. M. Wonham. Decoupling and pole assignment by dynamic compensation. SIAM Journal on Control, 8(3):317–337, August 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Narendra:1968:SAN

[NC68] Kumpati S. Narendra and Yo-Sung Cho. Stability analysis of nonlinear and time-varying discrete systems. SIAM Journal on Control, 6(4):625– 646, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Neustadt:1962:MEC

[Neu62] Lucien W. Neustadt. Minimum effort control systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1 (1):16–31, ???? 1962. CODEN JSM-CAL. ISSN 0887-4603.

Neustadt:1964:OMP

[Neu64] Lucien W. Neustadt. Optimization, a moment problem, and nonlinear programming. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1): 33–53, ???? 1964. CODEN JSM-CAL. ISSN 0887-4603.

Neustadt:1965:GTM

[Neu65] Lucien W. Neustadt. A general theory of minimum-fuel space trajectories. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2):317–356, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Neustadt:1966:AVT

[Neu66] Lucien W. Neustadt. An abstract variational theory with applications to a broad class of optimization problems. I. General theory. SIAM Journal on Control, 4(3):505–527, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Neustadt:1967:AVT

[Neu67] Lucien W. Neustadt. An abstract variational theory with applications to a broad class of optimization problems. II. Applications. SIAM Journal on Control, 5(1):90–137, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Nishiura:1967:ETO

[Nis67] Togo Nishiura. On an existence theorem for optimal control. SIAM Journal on Control, 5(4):532–544, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Narendra:1966:SCD

[NN66] Kumpati S. Narendra and Charles P.

Neuman. Stability of a class of differential equations with a single monotone nonlinearity. SIAM Journal on Control, 4(2):295–308, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Noldus:1975:ICT

[Nol75] E. Noldus. Instability criteria for time-varying nonlinear functional differential systems. SIAM Journal on Control, 13(2):420–433, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Norris:1967:LSP

[Nor67] D. O. Norris. Lagrangian saddle points and optimal control. SIAM Journal on Control, 5(4):594–599, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Nor68].

Norris:1968:ELS

[Nor68] D. O. Norris. Erratum: Lagrangian saddle points and optimal control. SIAM Journal on Control, 6(4):647, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Nor67].

Norris:1971:GLM

[Nor71] D. O. Norris. A generalized Lagrange multiplier rule for equality constraints in normed linear spaces. SIAM Journal on Control, 9(4):561– 567, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Norris:1975:CLS

[NS75] D. O. Norris and L. E. Snyder. Consistency of least-squares estimates used in linear systems identification. SIAM Journal on Control, 13(6): 1183–1193, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Neustadt:1970:CPO

[NW70] L. W. Neustadt and J. Warga. Comments on the paper “Optimal control of processes described by integral equations. I” by V. R. Vinokurov. SIAM Journal on Control, 8(4):572, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Vin69a].

Nelson:1968:CNO

[NY68] Paul Nelson, Jr. and Gale Young. A class of nonstandard optimal control problems with application to nuclear reactor economics. SIAM Journal on Control, 6(2):281–302, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Oguztoreli:1963:RTC

[Oğu63a] M. N. Oğuztörelí. Relay type control systems with retardation and switching delay. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3): 275–289, ???? 1963. CODEN JSM-CAL. ISSN 0887-4603.

Oguztoreli:1963:TOC

[Oğu63b] M. N. Oğuztörelí. A time optimal control problem for systems described by differential difference equations. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3):290–310, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Oguztoreli:1964:OPS

[Ogu64] M. Namík Oguztöreli. Optimal pursuit strategy processes with retarded control systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1): 89–105, ???? 1964. CODEN JSM-CAL. ISSN 0887-4603.

Oza:1968:SIP

[OJ68] K. G. Oza and E. I. Jury. System identification and the principle of random contraction mapping. SIAM Journal on Control, 6(2):244– 257, ???? 1968. CODEN SJCOA9.

ISSN 0036-1402 (print), 2469-4231 (electronic).

OMalley:1975:SPL

[OK75] R. E. O’Malley, Jr. and C. F. Kung. The singularly perturbed linear state regulator problem. II. SIAM Journal on Control, 13(2):327–337, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Okamura:1964:SMT

[Oka64] Kiyohisa Okamura. Some mathematical theory of the penalty method for solving optimum control problems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):317–331, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Olech:1974:CWC

[Ole74] Czeslaw Olech. Characterization of the weak closure of certain sets of integrable functions. SIAM Journal on Control, 12(2):311–318, May 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

OMalley:1972:SPL

[O’M72] R. E. O’Malley, Jr. The singularly perturbed linear state regulator problem. SIAM Journal on Control, 10(3):399–413, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Onuchic:1971:IPT

[Onu71] Nelson Onuchic. Invariance properties in the theory of ordinary differential equations with applications to stability problems. SIAM Journal on Control, 9(1):97–104, ???? 1971.

CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

PalludelaBarriere:1966:DDO

[Pal66] R. Pallu de la Barrière. Duality in dynamic optimization. SIAM Journal on Control, 4(1):159–163, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Papoulis:1973:DAA

[Pap73] A. Papoulis. A different approach to the analysis of tracer data. SIAM Journal on Control, 11(3):466–474, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Parzen:1962:EDP

[Par62] Emanuel Parzen. Extraction and detection problems and reproducing kernel Hilbert spaces. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1 (1):35–62, ???? 1962. CODEN JSM-CAL. ISSN 0887-4603.

Pau:1975:DGN

[Pau75] L. F. Pau. Differential games and a Nash equilibrium searching algorithm. SIAM Journal on Control, 13(4):835–852, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Paz:1967:FSS

[Paz67] A. Paz. A finite set of n × n stochastic matrices generating all n-dimensional probability vectors whose coordinates have finite binary expansion. SIAM Journal on Control, 5(4):545–554, ???? 1967. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Potter:1970:MBE

[PD70] James E. Potter and James C. Deckert. The minimal bound on the estimation error covariance matrix in the presence of correlated driving noise. SIAM Journal on Control, 8 (4):513–526, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pearson:1966:DDT

[Pea66a] J. D. Pearson. Duality and a decomposition technique. SIAM Journal on Control, 4(1):164–172, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pearson:1966:DBE

[Pea66b] J. D. Pearson. On the duality between estimation and control. SIAM Journal on Control, 4(4):594–600, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Peleg:1972:EPO

[Pel72] Bezalel Peleg. Efficiency prices for optimal consumption plans. IV. SIAM Journal on Control, 10(3): 414–433, ???? 1972. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Peng:1972:ISB

[Pen72] T. K. C. Peng. Invariance and stability for bounded uncertain systems. SIAM Journal on Control, 10(4): 679–690, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Petrosyan:1972:TGT

[Pet72] L. A. Petrosyan. Topological games and their applications to pursuit problems. I. SIAM Journal on Control, 10(1):194–202, ???? 1972. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Planchart:1975:EAS

[PH75] Alejo Planchart and Arthur P.

Hurter, Jr. Efficient algorithm for the solution of the Weber problem with mixed norms. SIAM Journal on Control, 13(3):650–665, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pliska:1973:MCD

[Pli73] Stanley R. Pliska. Multiperson controlled diffusions. SIAM Journal on Control, 11(4):563–586, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pliska:1975:SRM

[Pli75] Stanley R. Pliska. Semigroup representation of the maximum expected reward vector in continuous parameter Markov decision theory. SIAM Journal on Control, 13(6): 1115–1129, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pecsvaradi:1970:NIP

[PN70] Thomas Pecsvaradi and Kumpati S.

Narendra. New iterative procedure for the minimization of a quadratic form on a convex set. SIAM Journal on Control, 8(3):396–402, August 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pontryagin:1965:SDG

[Pon65] L. S. Pontryagin. On some differential games. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1): 49–52, ???? 1965. CODEN JSM-CAL. ISSN 0887-4603.

Ponstein:1968:MFO

[Pon68] J. Ponstein. Multiplier functions in optimal control. SIAM Journal on Control, 6(4):648–658, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pontryagin:1974:LDG

[Pon74] L. S. Pontryagin. Linear differential games. SIAM Journal on Control, 12(2):262–267, May 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Popov:1972:IDL

[Pop72] V. M. Popov. Invariant description of linear, time-invariant controllable systems. SIAM Journal on Control, 10(2):252–264, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Porter:1966:OCD

[Por66] William A. Porter. On the optimal control of distributive systems. SIAM Journal on Control, 4(3):466– 472, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Porter:1967:FSP

[Por67] William A. Porter. On function space pursuit-evasion games. SIAM Journal on Control, 5(4):555–574, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Porter:1968:MPA

[Por68] William A. Porter. A minimization problem and its applications to optimal control and system sensitivity. SIAM Journal on Control, 6(2):303– 311, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pironneau:1973:DMO

[PP73] O. Pironneau and E. Polak. A dual method for optimal control problems with initial and final boundary constraints. SIAM Journal on Control, 11(3):534–549, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Parthasarathy:1975:ESP

[PR75] T. Parthasarathy and T. E. S.

Raghavan. Existence of saddle points and Nash equilibrium points for differential games. SIAM Journal on Control, 13(5):977–980, August 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Prekopa:1966:PDO

[Prè66] András Prèkopa. On the probability distribution of the optimum of a random linear program. SIAM Journal on Control, 4(1):211–222, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Pshenichniy:1966:LOC

[Psh66] B. N. Pshenichniy. Linear optimal control problems. SIAM Journal on Control, 4(4):577–593, ???? 1966.

CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rao:1972:CDG

[Rao72] A. N. V. Rao. Comparison of differential games of fixed duration. SIAM Journal on Control, 10(2):393–397, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rao:1974:NSI

[RBR74] N. J. Rao, J. D. Borwanker, and D. Ramkrishna. Numerical solution of Ito integral equations. SIAM Journal on Control, 12(1):124–139, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Ross:1969:OCP

[RFL69] D. W. Ross and I. Flügge-Lotz. An optimal control problem for systems with differential-difference equation dynamics. SIAM Journal on Control, 7(4):609–623, November 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rishel:1965:EPP

[Ris65] Raymond W. Rishel. An extended Pontryagin principle for control systems whose control laws contain measures. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2):191–205, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Rishel:1966:ECR

[Ris66] R. W. Rishel. An example concerning rockets capable of impulsive thrust. SIAM Journal on Control, 4(4):740–744, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rishel:1970:NSD

[Ris70] Raymond Rishel. Necessary and sufficient dynamic programming conditions for continuous time stochastic optimal control. SIAM Journal on Control, 8(4):559–571, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rishel:1971:WSP

[Ris71a] Raymond W. Rishel. Weak solutions of a partial differential equation of dynamic programming. SIAM Journal on Control, 9(4):519–528, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rissanen:1971:RIL

[Ris71b] J. Rissanen. Recursive identification of linear systems. SIAM Journal on Control, 9(3):420–430, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rishel:1975:DPM

[Ris75] Raymond Rishel. Dynamic programming and minimum principles for systems with jump Markov disturbances. SIAM Journal on Control, 13(2):338–371, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rink:1968:CCB

[RM68] R. E. Rink and R. R. Mohler. Completely controllable bilinear systems. SIAM Journal on Control, 6(3):477– 486, ???? 1968. CODEN SJCOA9.

ISSN 0036-1402 (print), 2469-4231 (electronic).

Robinson:1973:LSM

[RM73] Stephen M. Robinson and Robert R. Meyer. Lower semicontinuity of multivalued linearization mappings. SIAM Journal on Control, 11(3): 525–533, ???? 1973. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Robinson:1975:AEB

[Rob75] Stephen M. Robinson. Application of error bounds for convex programming in a linear space. SIAM Journal on Control, 13(2):271–273, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rockafellar:1972:SCC

[Roc72] R. Tyrrell Rockafellar. State constraints in convex control problems of Bolza. SIAM Journal on Control, 10(4):691–715, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rockafellar:1974:ALM

[Roc74] R. Tyrrell Rockafellar. Augmented Lagrange multiplier functions and duality in nonconvex programming. SIAM Journal on Control, 12(2): 268–285, May 1974. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Roitenberg:1970:ONS

[Roĭ70] Ye. Ya. Roĭtenberg. Observability of nonlinear systems. SIAM Journal on Control, 8(3):338–345, ???? 1970.

CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Romeo:1969:SBF

[Rom69] Frank P. Romeo, Jr. Sequentially best filter. SIAM Journal on Control, 7(4):596–608, ???? 1969. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Root:1969:OCN

[Roo69] James G. Root. Optimum control of non-Gaussian linear stochastic systems with inaccessible state variables. SIAM Journal on Control, 7(2):317–323, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Root:1975:MSIa

[Roo75a] William L. Root. On the modeling of systems for identification. i: ε-representations of classes of systems. SIAM Journal on Control, 13(4):927–944, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Root:1975:MSIb

[Roo75b] William L. Root. On the modeling of systems for identification. Part II: Time-varying systems. SIAM Journal on Control, 13(4):945–974, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rosenbrock:1964:SCS

[Ros64] H. H. Rosenbrock. Some conditions for the stability of nonlinear time-dependent differential equations. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):171–180, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Rosen:1966:ISN

[Ros66] J. B. Rosen. Iterative solution of nonlinear optimal control problems. SIAM Journal on Control, 4(1):223– 244, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Roxin:1965:SCS

[Rox65] Emilio Roxin. On stability in control systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3):357–372, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Ramarajan:1973:CSM

[RT73] S. Ramarajan and M. A. L. Thathachar. Construction of stability multipliers with prescribed phase characteristics: an improved value for σ. SIAM Journal on Control, 11(1):94–99, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rubio:1966:FPM

[Rub66] J. E. Rubio. A fixed-point method for a minimum-norm control problem. SIAM Journal on Control, 4 (4):705–715, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rublein:1972:PCC

[Rub72] G. T. Rublein. On pursuit with curvature constraints. SIAM Journal on Control, 10(1):37–39, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Rubio:1975:GCE

[Rub75] J. E. Rubio. Generalized curves and extremal points. SIAM Journal on Control, 13(1):28–47, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Russell:1964:PFB

[Rus64] D. L. Russell. Penalty functions and bounded phase coordinate control. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):409–422, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Russell:1966:ORL

[Rus66] David L. Russell. Optimal regulation of linear symmetric hyperbolic systems with finite dimensional controls. SIAM Journal on Control, 4 (2):276–294, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Russell:1969:SPP

[Rus69a] C. Bradley Russell. Surveillance problems. Poisson process under costly surveillance. SIAM Journal on Control, 7(4):624–636, November 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Russell:1969:CST

[Rus69b] David L. Russell. Continuity in the strong topology of operatorvalued solutions of nonlinear differential equations with an application to optimal control. SIAM Journal on Control, 7(1):132–140, ???? 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Russell:1971:BVCa

[Rus71a] David L. Russell. Boundary value control of the higher-dimensional wave equation. SIAM Journal on Control, 9(1):29–42, February 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Russell:1971:BVCb

[Rus71b] David L. Russell. Boundary value control theory of the higherdimensional wave equation,Part II. SIAM Journal on Control, 9(3):401– 419, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Russell:1973:QPC

[Rus73] David L. Russell. Quadratic performance criteria in boundary control of linear symmetric hyperbolic systems. SIAM Journal on Control, 11 (3):475–509, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Russak:1975:SON

[Rus75] I. Bert Russak. Second order necessary conditions for problems with state inequality constraints. SIAM Journal on Control, 13(2):372–388, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Silverman:1968:COS

[SA68] L. M. Silverman and B. D. O. Anderson. Controllability, observability and stability of linear systems. SIAM Journal on Control, 6(1):121– 130, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sakawa:1966:SOP

[Sak66] Yoshiyuki Sakawa. On a solution of an optimization problem in linear systems with quadratic performance index. SIAM Journal on Control, 4 (2):382–395, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sakawa:1970:SLP

[Sak70] Yoshiyuki Sakawa. Solution of linear pursuit-evasion games. SIAM Journal on Control, 8(1):100–112 (or 101–112??), February 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sakawa:1974:CPD

[Sak74] Yoshiyuki Sakawa. Controllability for partial differential equations of parabolic type. SIAM Journal on Control, 12(3):389–400, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sakawa:1975:ORP

[Sak75] Yoshiyuki Sakawa. Observability and related problems for partial differential equations of parabolic type. SIAM Journal on Control, 13(1):14– 27, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sandberg:1964:SSS

[San64] I. W. Sandberg. On the solutions of systems of second order differential equations with variable coefficients. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):192–198, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Sannuti:1975:AES

[San75] Peddapullaiah Sannuti. Asymptotic expansions of singularly perturbed quasi-linear optimal systems. SIAM Journal on Control, 13(3):572–592, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Saperstone:1973:GCL

[Sap73] Stephen H. Saperstone. Global controllability of linear systems with positive controls. SIAM Journal on Control, 11(3):417–423, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sasai:1972:NPM

[Sas72] Hitoshi Sasai. Note on the penalty method for distributed parameter optimal control problems. SIAM Journal on Control, 10(4):730–736, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sasai:1974:IPM

[Sas74] Hitoshi Sasai. Interior penalty method for minimax problems with constraints. SIAM Journal on Control, 12(4):643–649, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Steding:1974:INI

[SB74] Thomas L. Steding and Arthur R. Bergen. Instability of nonlinear infinite-dimensional feedback systems using Lyapunov functionals. SIAM Journal on Control, 12(4): 664–678, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Scalzo:1974:DGR

[Sca74a] Richard C. Scalzo. Differential games with restricted phase coordinates. SIAM Journal on Control, 12 (3):426–435, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Scalzo:1974:PLQ

[Sca74b] Richard C. Scalzo. N-person linearquadratic differential games with constraints. SIAM Journal on Control, 12(3):419–425, ???? 1974. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Schmaedeke:1965:OCT

[Sch65] W. W. Schmaedeke. Optimal control theory for nonlinear vector differential equations containing measures. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(2):231–280, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Schmaedeke:1967:MTO

[Sch67] Wayne Schmaedeke. Mathematical theory of optimal control for semilinear hyperbolic systems in two independent variables. SIAM Journal on Control, 5(1):138–152, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Schwarzkopf:1972:OCP

[Sch72] A. B. Schwarzkopf. Optimal controls for problems with a restricted state space. SIAM Journal on Control, 10 (3):487–511, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Schultz:1973:KTA

[Sch73] Hilbert K. Schultz. A Kuhn–Tucker algorithm. SIAM Journal on Control, 11(3):438–445, ???? 1973. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Schmookler:1975:SMS

[Sch75a] Yulia Schmookler. Statistical mechanics and systems of large numbers of elements with random interaction. SIAM Journal on Control, 13(2):243–270, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Schwarzkopf:1975:CTN

[Sch75b] A. B. Schwarzkopf. Controllability and tenability of nonlinear systems with state equality constraints. SIAM Journal on Control, 13(3): 695–705, May 1975. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Schwarzkopf:1975:RCP

[Sch75c] A. B. Schwarzkopf. Relaxed control problems with state equality constraints. SIAM Journal on Control, 13(3):677–694, May 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Soults:1971:DGV

[SD71] Donald J. Soults and H. T. David. The distribution of 2×n game values and program optima. SIAM Journal on Control, 9(1):135–141, February 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Szego:1963:RNP

[SG63] G. P. Szegő and G. R. Geiss. A remark on “A New Partial Differential Equation for the Stability Analysis of Time Invariant Control Systems”. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(3):369–376, ???? 1963. CODEN JSMCAL. ISSN 0887-4603.

Sinha:1971:SND

[SH71] A. S. C. Sinha and R. G. Hoft.

Stability of a nonautonomous differential equation of fourth order. SIAM Journal on Control, 9(1):8–14, February 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Skoog:1972:DSR

[Sko72] R. A. Skoog. On dynamical systems realizing stationary weighting patterns and time-varying feedback systems. SIAM Journal on Control, 10(1):48–55, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Skoog:1974:PCI

[Sko74] Ronald A. Skoog. Positivity conditions and instability criteria for feedback systems. SIAM Journal on Control, 12(1):83–98, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Shafran:1970:RPB

[SL70] J. S. Shafran and J. Y. S. Luh.

Recoverability for processes with bounded control amplitudes and rates. SIAM Journal on Control, 8(2):189–201, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Slemrod:1974:NCC

[Sle74] Marshall Slemrod. A note on complete controllability and stabilizability for linear control systems in Hilbert space. SIAM Journal on Control, 12(3):500–508, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Silverman:1967:COT

[SM67] L. M. Silverman and H. E. Meadows. Controllability and observability in time-variable linear systems. SIAM Journal on Control, 5(1):64–73, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Smith:1964:MLF

[Smi64] Gerald L. Smith. Multivariable linear filter theory applied to space vehicle guidance. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(1): 19–32, ???? 1964. CODEN JSM-CAL. ISSN 0887-4603.

Snow:1964:SOC

[Sno64] Donald R. Snow. Singular optimal controls for a class of minimum effort problems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):203–219, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Snyders:1975:EEO

[Sny75] Jakov Snyders. Error expressions for optimal stationary state estimation. SIAM Journal on Control, 13(5): 1093–1102, August 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sarma:1971:NNC

[SP71a] I. G. Sarma and U. R. Prasad.

A note on the necessary conditions for optimal strategies in a class of noncooperative N-person differential games. SIAM Journal on Control, 9(3):441–445, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Silverman:1971:IOS

[SP71b] L. M. Silverman and H. J. Payne.

Input-output structure of linear systems with application to the decoupling problem. SIAM Journal on Control, 9(2):199–233, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Schmaedeke:1964:TOC

[SR64] W. W. Schmaedeke and D. L. Russell. Time optimal control with amplitude and rate limited controls. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):373–395, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Sarma:1969:NCO

[SRP69] I. G. Sarma, R. K. Ragade, and U. R. Prasad. Necessary conditions for optimal strategies in a class of noncooperative N-person differential games. SIAM Journal on Control, 7(4):637–644, November 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sasai:1971:CAS

[SS71] Hitoshi Sasai and Etsujro Shimemura. On the convergence of approximating solutions for linear distributed parameter optimal control problems. SIAM Journal on Control, 9(2):263– 273, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Stern:1972:ABV

[Ste72] Ronald J. Stern. Asymptotic behavior of the value of differential games. SIAM Journal on Control, 10(4): 747–753, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Stern:1973:OLA

[Ste73] Ronald J. Stern. Open loop approximation of differential games. SIAM Journal on Control, 11(1):20– 31, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Stern:1974:DGS

[Ste74] Ronald J. Stern. Differential games of survival with space-like terminal set. SIAM Journal on Control, 12(1): 167–177, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Stone:1973:NSC

[Sto73] Lawrence D. Stone. Necessary and sufficient conditions for optimal control of semi-Markov jump processes. SIAM Journal on Control, 11(2): 187–201, ???? 1973. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Stratonovich:1966:NRS

[Str66] R. L. Stratonovich. A new representation for stochastic integrals and equations. SIAM Journal on Control, 4(2):362–371, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Suryanarayana:1973:NCO

[Sur73] M. B. Suryanarayana. Necessary conditions for optimization problems with hyperbolic partial differential equations. SIAM Journal on Control, 11(1):130–147, ???? 1973. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sussmann:1972:BBP

[Sus72] Hector J. Sussmann. The “bang– bang” problem for certain control systems in GL(n, R). SIAM Journal on Control, 10(3):470–476, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sussmann:1975:NDN

[Sus75] Héctor J. Sussmann. On the number of directions needed to achieve controllability. SIAM Journal on Control, 13(2):414–419, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sonneborn:1964:BBP

[SV64] L. M. Sonneborn and F. S. Van Vleck. The bang–bang principle for linear control systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(2):151–159, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Snyders:1975:RLS

[SW75] J. Snyders and W. M. Wonham. Regulation of linear stochastic systems. SIAM Journal on Control, 13(4):853–864, July 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Sworder:1964:MCD

[Swo64] D. D. Sworder. Minimax control of discrete time stochastic systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):433–449, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Saperstone:1971:CLO

[SY71] Stephen H. Saperstone and James A. Yorke. Controllability of linear oscillatory systems using positive controls. SIAM Journal on Control, 9 (2):253–262, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Szego:1962:NPD

[Sze62] G. P. Szegő. On a new partial differential equation for the stability analysis of time invariant control systems. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 1(1): 63–75, ???? 1962. CODEN JSM-CAL. ISSN 0887-4603.

Tokumaru:1970:DMM

[TAG70] H. Tokumaru, N. Adachi, and K. Goto. Davidon’s method for minimization problems in Hilbert space with an application to control problems. SIAM Journal on Control, 8(2):163–178, May 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Chen:1968:SLT

[tC68] Chi tsong Chen. L_{p}-stability of linear time-varying feedback systems. SIAM Journal on Control, 6(2):186– 193, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Triggiani:1975:COB

[Tri75] Roberto Triggiani. Controllability and observability in Banach space with bounded operators. SIAM Journal on Control, 13(2):462–491, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Troch:1971:NCC

[Tro71] Inge Troch. A note on complete controllability. SIAM Journal on Control, 9(4):543–546, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Tsujioka:1970:RCS

[Tsu70] Kunio Tsujioka. Remarks on controllability of second order evolution equations in Hilbert spaces. SIAM Journal on Control, 8(1):90– 99, February 1970. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Topkis:1967:CSF

[TV67] Donald M. Topkis and Arthur F.

Veinott, Jr. On the convergence of some feasible direction algorithms for nonlinear programming. SIAM Journal on Control, 5(2):268–279, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Thompson:1975:LQC

[TV75] D. D. Thompson and R. A. Volz. Linear quadratic cost problem with linear state constraints and the nonsymmetric Riccati equation. SIAM Journal on Control, 13(1):110–145, January 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Vachino:1966:SDI

[Vac66] Rinaldo F. Vachino. Steepest descent with inequality constraints on the control variables. SIAM Journal on Control, 4(1):245–261, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Varaiya:1966:DLS

[Var66] P. Varaiya. Decomposition of largescale systems. SIAM Journal on Control, 4(1):173–178, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Varaiya:1967:ESD

[Var67] P. P. Varaiya. On the existence of solutions to a differential game. SIAM Journal on Control, 5(1):153– 162, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Varaiya:1970:PNS

[Var70] Pravin Varaiya. N-person nonzero sum differential games with linear dynamics. SIAM Journal on Control, 8(4):441–449, ???? 1970. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Venkatesh:1971:GVC

[Ven71] Y. V. Venkatesh. Global variation criteria for stability of linear time varying systems. SIAM Journal on Control, 9(3):431–440, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Vidyasagar:1972:IOS

[Vid72] Mathukumalli Vidyasagar. Inputoutput stability of a broad class of linear time-invariant multivariable systems. SIAM Journal on Control, 10(1):203–209, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Vidyasagar:1975:CFS

[Vid75] Mathukumalli Vidyasagar. Coprime factorizations and stability of multivariable distributed feedback systems. SIAM Journal on Control, 13 (6):1144–1155, November 1975. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Vinokurov:1969:OCPa

[Vin69a] V. R. Vinokurov. Optimal control of processes described by integral equations. I. SIAM Journal on Control, 7(2):324–336, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Vin70] and comments [NW70].

Vinokurov:1969:OCPb

[Vin69b] V. R. Vinokurov. Optimal control of processes described by integral equations. II. SIAM Journal on Control, 7(2):337–345, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Vinokurov:1969:OCPc

[Vin69c] V. R. Vinokurov. Optimal control of processes described by integral equations. III. SIAM Journal on Control, 7(2):346–355, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Vinokurov:1970:EOC

[Vin70] V. R. Vinokurov. Erratum: “Optimal control of processes described by integral equations. I”. SIAM Journal on Control, 8(3):440, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Vin69a].

Vinokurov:1973:FCP

[Vin73] V. R. Vinokurov. Further comments on the paper: “Optimal control of processes described by integral equations. I”. (Izv. Vysš. Učebn. Zaved. Matematika 62, no. 7 (1976), 21–33) by V. R. Vinokurov. SIAM Journal on Control, 11(1):185, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Vinter:1974:GDB

[Vin74] Richard B. Vinter. Generalization to dual Banach spaces of a theorem by Balakrishnan. SIAM Journal on Control, 12(1):150–166, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Virsan:1970:NCO

[Vır70] C. Vırsan. Necessary conditions for optimization problems with operatorial constraints. SIAM Journal on Control, 8(4):527–558, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Varaiya:1966:BIB

[VL66] P. P. Varaiya and R. Liu. Boundedinput bounded-output stability of nonlinear time-varying differential systems. SIAM Journal on Control, 4(4):698–704, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Varaiya:1969:ESP

[VL69] P. Varaiya and Jiguan Lin. Existence of saddle points in differential games. SIAM Journal on Control, 7(1):141–157, February 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

VanSlyke:1966:PUU

[VW66] Richard Van Slyke and Roger Wets.

Programming under uncertainty and stochastic optimal control. SIAM Journal on Control, 4(1):179–193, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wahba:1973:MQF

[Wah73] Grace Wahba. On the minimization of a quadratic functional subject to a continuous family of linear inequality constraints. SIAM Journal on Control, 11(1):64–79, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wang:1975:OCP

[Wan75] P. K. C. Wang. Optimal control of parabolic systems with boundary conditions involving time delays. SIAM Journal on Control, 13(2): 274–293, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Warga:1965:MPU

[War65a] J. Warga. Minimax problems with unilateral curves in the calculus of variations. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(1):91–105, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Warga:1965:VPU

[War65b] J. Warga. Variational problems with unbounded controls. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 3(3):424–438, ???? 1965. CODEN JSMCAL. ISSN 0887-4603.

Warga:1967:FRC

[War67a] J. Warga. Functions of relaxed controls. SIAM Journal on Control, 5 (4):628–641, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Warga:1967:RMF

[War67b] J. Warga. Restricted minima of functions of controls. SIAM Journal on Control, 5(4):642–656, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Warga:1970:CPF

[War70a] J. Warga. Control problems with functional restrictions. SIAM Journal on Control, 8(3):360–371, August 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Warga:1970:UMC

[War70b] J. Warga. Unilateral and minimax control problems defined by integral equations. SIAM Journal on Control, 8(3):372–382, August 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Warga:1974:OCP

[War74] J. Warga. Optimal controls with pseudodelays. SIAM Journal on Control, 12(2):286–299, May 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wu:1969:SNT

[WD69] M. Y. Wu and C. A. Desoer. L^{p} stability (1 ≤ p ≤ ∞) of nonlinear time-varying feedback systems. SIAM Journal on Control, 7(2):356– 364, May 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Warren:1975:DCS

[WE75] Michael E. Warren and Adrian E.

Eckberg, Jr. On the dimensions of controllability subspaces: a characterization via polynomial matrices and Kronecker invariants. SIAM Journal on Control, 13(2):434–445, February 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Weiss:1967:CDD

[Wei67] Leonard Weiss. On the controllability of delay-differential systems. SIAM Journal on Control, 5(4):575– 587, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Weiss:1968:STL

[Wei68] Leonard Weiss. On the structure theory of linear differential systems.

SIAM Journal on Control, 6(4):659– 680, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Weiss:1972:CRS

[Wei72] Leonard Weiss. Controllability, realization and stability of discrete-time systems. SIAM Journal on Control, 10(2):230–251, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Weiner:1977:CGP

[Wei77] L. B. Weiner. Comments on: “Generalized Prediction-Correction Estimation”. SIAM Journal on Control and Optimization, 15(5):857– 858, August 1977. CODEN SJ-CODC. ISSN 0363-0129 (print), 1095-7138 (electronic). See [KH69].

Wolovich:1969:SMS

[WF69] W. A. Wolovich and P. L. Falb. On the structure of multivariable systems. SIAM Journal on Control, 7 (3):437–451, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wierzbicki:1972:MPS

[Wie72] Andrzej P. Wierzbicki. Maximum principle for semiconvex performance functionals. SIAM Journal on Control, 10(3):444–459, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Willems:1969:SII

[Wil69] Jan C. Willems. Stability, instability, invertibility and causality. SIAM Journal on Control, 7(4):645–671, November 1969. CODEN SJCOA9.

ISSN 0036-1402 (print), 2469-4231 (electronic).

Willems:1971:GLF

[Wil71a] Jan C. Willems. The generation of Lyapunov functions for inputoutput stable systems. SIAM Journal on Control, 9(1):105–134, February 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Willman:1971:SEM

[Wil71b] W. W. Willman. Some effects of measurement uncertainty in linear multistage games. SIAM Journal on Control, 9(2):274–286, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Witsenhausen:1968:CSO

[Wit68] H. S. Witsenhausen. A counterexample in stochastic optimum control. SIAM Journal on Control, 6(1):131– 147, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Witsenhausen:1970:PBU

[Wit70] H. S. Witsenhausen. On performance bounds for uncertain systems. SIAM Journal on Control, 8(1):55– 89, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Witsenhausen:1971:ISF

[Wit71] H. S. Witsenhausen. On information structures, feedback and causality. SIAM Journal on Control, 9 (2):149–160, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wonham:1970:DPA

[WM70] W. M. Wonham and A. S. Morse. Decoupling and pole assignment in linear multivariable systems: a geometric approach. SIAM Journal on Control, 8(1):1–18, February 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wolovich:1972:USF

[Wol72] W. A. Wolovich. The use of state feedback for exact model matching. SIAM Journal on Control, 10 (3):512–523, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wonham:1964:SAS

[Won64] W. M. Wonham. Some applications of stochastic differential equations to optimal nonlinear filtering. Journal of the Society for Industrial and Applied Mathematics. Series A, On control, 2(3):347–369, ???? 1964. CODEN JSMCAL. ISSN 0887-4603.

Wonham:1967:OSC

[Won67] W. M. Wonham. Optimal stationary control of a linear system with state-dependent noise. SIAM Journal on Control, 5(3):486–500, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wonham:1968:MRE

[Won68a] W. M. Wonham. On a matrix Riccati equation of stochastic control. SIAM Journal on Control, 6(4):681– 697, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See erratum [Won69].

Wonham:1968:STS

[Won68b] W. M. Wonham. On the separation theorem of stochastic control. SIAM Journal on Control, 6(2):312– 326, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wonham:1969:EMR

[Won69] W. M. Wonham. Erratum: “On a matrix Riccati equation of stochastic control”. SIAM Journal on Control, 7(2):365, ???? 1969. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic). See [Won68a].

Wong:1971:RMQ

[Won71] E. Wong. Representation of martingales, quadratic variation and applications. SIAM Journal on Control, 9(4):621–633, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wonham:1973:TRL

[Won73] W. M. Wonham. Tracking and regulation in linear multivariable systems. SIAM Journal on Control, 11 (3):424–437, ???? 1973. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wonham:1974:RIS

[WP74] W. M. Wonham and J. B. Pearson. Regulation and internal stabilization in linear multivariable systems. SIAM Journal on Control, 12 (1):5–18, February 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wagner:1974:NER

[WS74a] Daniel H. Wagner and Lawrence D. Stone. Necessity and existence results on constrained optimization of separable functionals by a multiplier rule. SIAM Journal on Control, 12 (3):356–372, ???? 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Wagner:1974:OAU

[WS74b] Daniel H. Wagner and Lawrence D. Stone. Optimization of allocations under a coverability condition. SIAM Journal on Control, 12(3): 373–379, ???? 1974. CODEN SJ-COA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Walkup:1969:SPR

[WW69] David W. Walkup and Roger J.- B. B. Wets. Some practical regularity conditions for nonlinear programs. SIAM Journal on Control, 7 (3):430–436, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Yorke:1971:APL

[Yor71] James A. Yorke. Another proof of the Liapunov convexity theorem. SIAM Journal on Control, 9(3):351– 353, ???? 1971. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Yorke:1972:MPC

[Yor72] James A. Yorke. The maximum principle and controllability of nonlinear equations. SIAM Journal on Control, 10(2):334–338, ???? 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zabczyk:1974:RCD

[Zab74] J. Zabczyk. Remarks on the control of discrete-time distributed parameter systems. SIAM Journal on Control, 12(4):721–735, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zabczyk:1975:OSC

[Zab75] Jerzy Zabczyk. On optimal stochastic control of discrete-time systems in Hilbert space. SIAM Journal on Control, 13(6):1217–1234, November 1975. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zakai:1967:UBM

[Zak67] Moshe Zakai. On the ultimate boundedness of moments associated with solutions of stochastic differential equations. SIAM Journal on Control, 5(4):588–593, ???? 1967. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zakai:1969:LCE

[Zak69] Moshe Zakai. A Lyapunov criterion for the existence of stationary probability distributions for systems perturbed by noise. SIAM Journal on Control, 7(3):390–397, August 1969. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zemanian:1974:RHP

[Zem74] A. H. Zemanian. Relaxive Hilbert ports. SIAM Journal on Control, 12(1):106–123, February 1974. CO-DEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zames:1968:SCS

[ZF68] G. Zames and P. L. Falb. Stability conditions for systems with monotone and slope-restricted nonlinearities. SIAM Journal on Control, 6 (1):89–108, ???? 1968. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zlobec:1970:AKT

[Zlo70] Sanjo Zlobec. Asymptotic Kuhn– Tucker conditions for mathematical programming problems in a Banach space. SIAM Journal on Control, 8 (4):505–512, ???? 1970. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zmood:1974:ESC

[Zmo74] Ronald B. Zmood. Euclidean space controllability of control systems with delay. SIAM Journal on Control, 12(4):609–623, November 1974. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zolezzi:1972:NCO

[Zol72] Tullio Zolezzi. Necessary conditions for optimal controls of elliptic or parabolic problems. SIAM Journal on Control, 10(4):594–607, November 1972. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).

Zoutendijk:1966:NPN

[Zou66] G. Zoutendijk. Nonlinear programming: a numerical survey. SIAM Journal on Control, 4(1):194–210, ???? 1966. CODEN SJCOA9. ISSN 0036-1402 (print), 2469-4231 (electronic).